### v1

# One-Shot Real-World Demonstration Synthesis for Scalable Bimanual Manipulation

Abstract
Learning dexterous bimanual manipulation policies critically depends on large-scale, high-quality demonstrations, yet current paradigms face inherent trade-offs: teleoperation provides physically grounded data but is prohibitively labor-intensive, while simulation-based synthesis scales efficiently but suffers from sim-to-real gaps. We present BiDemoSyn, a framework that synthesizes contact-rich, physically feasible bimanual demonstrations from a single real-world example. The key idea is to decompose tasks into invariant coordination blocks and variable, object-dependent adjustments, then adapt them through vision-guided alignment and lightweight trajectory optimization. This enables the generation of thousands of diverse and feasible demonstrations within several hour, without repeated teleoperation or reliance on imperfect simulation. Across six dual-arm tasks, we show that policies trained on BiDemoSyn data generalize robustly to novel object poses and shapes, significantly outperforming recent baselines. By bridging the gap between efficiency and real-world fidelity, BiDemoSyn provides a scalable path toward practical imitation learning for complex bimanual manipulation without compromising physical grounding.
1 Introduction
Recent advances in robot manipulation have been propelled by data-driven imitation learning, where visuomotor policies trained on large-scale demonstration datasets enable dexterous single-arm and bimanual tasks previously deemed intractable. Methods like Action Chunking Transformer (ACT) zhao2023learning, Diffusion Policy (DP) chi2023diffusion and $\pi_{0}$ black2024pi0 exemplify this trend, achieving remarkable performance in contact-rich scenarios such as continuously pushing, battery assembly, and coordinated garment folding. These successes, however, hinge on a critical yet underappreciated premise: the availability of diverse, high-quality demonstrations, which are typically collected via expert teleoperation. As the community shifts toward more complex long-horizon bimanual tasks, the data scalability bottleneck becomes increasingly apparent. Even those largest manipulation datasets like RH20T fang2024rh20t, DROID khazatsky2024droid, Open X-Embodiment o2024open and AgiBot World bu2025agibot struggle to provide sufficient geometric and kinematic diversity for robust generalization, underscoring a pivotal question: how to scale real-world demonstration collection without compromising physical fidelity?
[…]
The answer, unfortunately, lies in a fundamental trade-off. Existing data acquisition methods fall into two paradigms, each with critical shortcomings. On one hand, human-operated systems (e.g., the ALOHA series zhao2023learning; fu2024mobile; zhao2024aloha) yield physically-accurate trajectories but demand prohibitive expertise, rendering them impractical for scaling to long-horizon tasks. On the other hand, simulation-based methods (e.g., MimicGen mandlekar2023mimicgen, RoboGen wang2024robogen, RoboCasa nasiriany2024robocasa, and DexMimicGen jiang2024dexmimicgen) bypass human labor through programmatic data generation utilizing massive 3D assets, yet discrepancies in contact dynamics and visual rendering inevitably plague real-world deployment bharadhwaj2024position; mccarthy2024towards. Thus, the tension between data quality (grounded in reality) and quantity (enabled by automation) persists as a key barrier to scalable imitation learning, particularly for contact-rich bimanual tasks requiring precise dual-arm coordination.
To bridge this gap, we present BiDemoSyn that synthesizes real-world bimanual demonstrations from a single exemplar, by algorithmically amplifying task semantics through a hierarchical decomposition of invariance and adaptability. Our key insight lies in recognizing that complex manipulation tasks exhibit a duality: while certain motion primitives (e.g., stable grasp sequences, dual-arm synchronization) remain invariant across instances, others (e.g., trajectory segments adapting to object shifting) require dynamic adjustments to contextual variations. Through a three-stage process, BiDemoSyn first deconstructs the initial demonstration into invariant primitives and geometry-sensitive components, then leverages visual perception to spatially align these elements with novel scenes, enabling generalization across object geometries and workspace layouts with minor human effort (primarily involving reinitializing manipulated objects). Finally, a physics-aware optimizer jointly modulates dual-arm trajectories, injecting diversity into adaptable components while enforcing real-world constraints on collision avoidance and kinematic coordination. This integrated approach achieves one-to-many demonstration synthesis: a single exemplar spawns hundreds of trajectories that preserve task intent while adapting to unseen configurations, all grounded in real-world dynamics rather than simulated proxies. Examples are shown in Fig. 1 and Fig. 2.
Unlike prior simulation-based methods mandlekar2023mimicgen; wang2024robogen; nasiriany2024robocasa; jiang2024dexmimicgen, BiDemoSyn operates entirely within the physical domain, ensuring synthesized data inherits the fidelity of human demonstrations. By unifying algorithmic scalability with physical realism, our framework empowers policies to transcend the limitations of narrow manually-collected data distributions. Experiments on real dual-arm platforms validate the effectiveness of BiDemoSyn. And visuomotor policies trained with our synthesized demonstrations can achieve superior success rates and generalization in tasks requiring precise contact coordination (e.g., small hole insertion, rotation angle control, purposeful handover and articulated object manipulation).
Our contributions are threefold: (1) One-Shot Synthesis Framework: A systematic pipeline combining task decomposition, vision-guided adaptation, and contact-aware trajectory optimization to generate scalable real-world bimanual demonstrations. (2) Reality-Grounded Data Generation: A completely simulator-free method for synthesizing bimanual demonstrations, ensuring physical fidelity by construction. (3) Empirical Validation in Complex Tasks: Comprehensive real-robot experiments demonstrating significant improvements in policy robustness and cross-configuration generalization on various bimanual manipulation tasks.
2 Related Works
Bimanual Robotic Manipulation. Prior literatures primarily target task-specific challenges (such as cloth folding colome2018dimensionality; weng2022fabricflownet; canberk2023cloth, untangling grannen2021untangling; peng2024tiebot, untwisting lin2024twisting; bahety2024screwmimic, and handover huang2023dynamic; li2023efficient) through handcrafted controllers or narrow skill libraries, limiting generalization due to rigid motion primitives and excessive assumptions. General methods xie2020deep; chitnis2020efficient; chen2022towards; chen2023bi; hartmann2022long; zhaodual2023afford either rely on predefined hierarchical arm roles (e.g., leader-follower krebs2022bimanual; liu2022robot; grotz2024peract2 and stabilizer-actor grannen2023stabilize; liu2024voxact) or hinge generalization entirely on dataset diversity, both of which are ultimately bottlenecked by the impractical costs of data acquisition. Recently, ALOHA series zhao2023learning; fu2024mobile; zhao2024aloha have revolutionized bimanual manipulation by dexterous low-cost teleoperation. Meanwhile, the visuomotor policy learning paradigm chi2023diffusion; ze2024dp3; yang2024equibot based on diffusion models ho2020denoising; song2021denoising has largely expanded the action modeling space and data utilization efficiency. Their followers team2024octo; kim2024openvla; liu2025rdt; black2024pi0; pertsch2025fast; lin2025data expect to train universal policies using large-scale teleoperation data but face scalability barriers from human effort. To improve reachability and dexterity, some studies use specialized hardware (e.g., multi-finger hands wang2024dexcap; shaw2024bimanual; fu2024humanplus; cheng2024open or tactile sensors lin2024learning; chen2024arcap), which complicate real-world adoption. In contrast, we utilize a fixed-base dual-arm platforrm with parallel grippers, synthesizing diverse demonstrations from a single example. By optimizing coordination as dynamic constraints rather than predefined hierarchies, we bypass hardware specificity and data scalability limitations, enabling task-agnostic policies grounded in physical feasibility.
Bimanual Demonstration Collection. The dominant path for acquiring bimanual demonstrations is human teleoperation khazatsky2024droid; bu2025agibot, which delivers high-quality data but suffers from prohibitive scalability costs. To alleviate it, two alternative routes have emerged: simulation-based synthesis garrett2024skillmimicgen; hua2024gensim2; liang2024make; yang2024physcene; wang2024gensim and learning from human videos ponimatkin20256d; ye2025video2policy; zhao2025taste; kareer2024egomimic; bharadhwaj2024track2act. The former, exemplified by MimicGen mandlekar2023mimicgen and DexMimicGen jiang2024dexmimicgen, augments a few demonstrations by generating variations in simulation using geometric transformations and kinematic constraints. Analogously, RoboGen wang2024robogen and RoboCasa nasiriany2024robocasa leverage 3D assets to procedurally generate full-scene manipulation data, yet such methods inevitably inherit sim-to-real gaps, from unrealistic textures to unphysical dynamics. The latter extracts bimanual hand-object manipulation from egocentric videos zhan2024oakink2; liu2024taco; grauman2024ego and maps them to dual arms via motion retargeting li2024okami; kerr2024robot; chen2024object; chen2024vividex or non-privileged representations (e.g., keypoints papagiannis2024rx; gao2024bi; wen2023any, affordances ju2024robo; nasiriany2024rt and correspondences peng2024tiebot; ko2024learning; zhang2024one). While human videos are abundant and low-cost, large morphological mismatches between humanity and robotic embodiments often need heuristic translation rules and hinder direct applicability. Our work navigates these trade-offs by synthesizing demonstrations directly in real-world. Given a single exemplar, we diversify trajectories through vision-based adaptation and physics-compliant optimization. This balances the fidelity and scalability, while avoiding retargeting hurdles, thereby enabling practical and scalable data acquisition for contact-rich bimanual tasks.
Several concurrent works exhibit notable limitations. DemoGen xue2025demogen and YOTO zhou2025you diversify demonstrations via 3D editing of the initial single-view point cloud, but perspective ambiguities induce visual artifacts in synthesized data. ODIL wang2025one avoids editing by segmenting objects for visual servoing and iteratively replanning trajectories, which is cumbersome for scalability. MoMaGen li2025momagen is a long-horizon mobile+bimanual demonstrations generator yet mainly focusing on the simulation domain. Our BiDemoSyn sidesteps these issues: synthesizing trajectories directly in real-world ensures visual-physical consistency without manual edits or robot replaying.
3 Preliminaries
Problem Formulation. Bimanual imitation learning aims to train visuomotor policies $\pi_{\theta}(\textbf{o}_{t})\rightarrow\textbf{a}_{t}$, where $\textbf{o}_{t}$ denotes multimodal observations (e.g., RGB images, joint states or point clouds) and $\textbf{a}_{t}$ represents dual-arm actions. While architectures like ACT zhao2023learning and DP chi2023diffusion enhance policy expressivity, their generalization critically depends on expert demonstrations $\mathcal{D}\!=\!\{\tau_{i}\}_{i=0}^{N}$ that densely span the feasible state-action manifold. Existing approaches either collect $\mathcal{D}$ via labor-intensive teleoperation or generate data in simulation, both failing to balance real-world fidelity and scalability. We address this gap by proposing the following problem:
Consider a bimanual task defined by an initial state $s_{0}$ and a specific goal $g$. Given a single real-world demonstration $\tau\!=\!\{\textbf{o}_{t},\textbf{a}_{t}\}_{t=0}^{T}$, our objective is to synthesize a dataset $\mathcal{D}_{syn}\!=\!\{\tau_{i}\}_{i=0}^{M}(M\gg 1)$ such that: (1) Task Consistency: Every $\tau_{i}\in\mathcal{D}_{syn}$ achieves the task goal $g$ under perturbed initial states (e.g., new object poses, scene layouts). (2) Physical Admissibility: Trajectories adhere to kinematic and dynamic constraints of the real-world system (e.g., collision avoidance, dual-arm coordination). (3) Diversity: $\mathcal{D}_{syn}$ covers task-relevant variations to enable robust policy training.
Visuomotor Policy Learning. Our $\pi_{\theta}$ maps observations $\textbf{o}_{t}$ to dual-arm actions $\textbf{a}_{t}\!=\!\left[\textbf{a}_{t}^{L},\textbf{a}_{t}^{R}\right]\!\in\!\mathbb{R}^{2d}$, where $\textbf{a}_{t}^{L}$, $\textbf{a}_{t}^{R}$ denotes left/right end-effector poses. We adapt a diffusion policy model chi2023diffusion; ze2024dp3; yang2024equibot to generate action sequences $\textbf{A}\!=\!\left[\textbf{a}_{t:{t+H}}^{L},\textbf{a}_{t:{t+H}}^{R}\right]\!\in\!\mathbb{R}^{2H\times d}$ over a horizon $H$. The diffusion process iteratively denoises a noisy action sequence $\textbf{A}_{k}$ toward kinematic feasibility over $K$ steps:
$\centering\textbf{A}_{k}=\sqrt{\alpha_{k}}\textbf{A}_{0}+\sqrt{1-\alpha_{k}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{\Sigma}),\@add@centering$ (1)
where $\alpha_{k}$ follows a cosine schedule nichol2021improved. And $\mathbf{\Sigma}\!\in\!\mathbb{R}^{2H\times 2H}$ is a block-diagonal covariance matrix $\left[\begin{array}[]{cc}\mathbf{\Sigma}_{L}&\rho\mathbf{\Sigma}_{LR}\\ \rho\mathbf{\Sigma}^{\top}_{LR}&\mathbf{\Sigma}_{R}\end{array}\right]$ encoding arm coordination priors, where $\mathbf{\Sigma}_{L}$, $\mathbf{\Sigma}_{R}$ governing temporal smoothness per arm, $\mathbf{\Sigma}_{LR}$ modeling inter-arm dependencies, and $\rho\!\in\![0,1]$ controlling coordination strength. Then, a denoiser $\mu_{\theta}$ processes the noised sequence conditioned on visual observations $\textbf{o}_{t}$ via:
$\centering\mu_{\theta}(\textbf{A}_{k},\textbf{o}_{t},k)=\texttt{UNet}_{\theta}(\texttt{Concat}\left[\textbf{A}_{k},\mathsf{E}(\textbf{o}_{t})\right],k),\@add@centering$ (2)
where $\mathsf{E}(\textbf{o}_{t})$ is a vision encoder (ResNet he2016deep or PointNet qi2017pointnet) extracting latent features. The $\texttt{UNet}_{\theta}$ employs asymmetric layers with early stages processing each arm independently, while deeper layers fusing bilateral coordinated features. Training minimizes:
$\centering\mathcal{L}(\theta)=\mathbb{E}_{\textbf{A}_{0},\textbf{o}_{t},k}\left[\left\|\mu_{\theta}(\textbf{A}_{k},\textbf{o}_{t},k)-\textbf{A}_{0}\right\|_{\textbf{W}}^{2}\right],\@add@centering$ (3)
where $\|\!\cdot\!\|_{\textbf{W}}$ is a Mahalanobis norm with $\textbf{W}\!=\!\mathbf{\Sigma}^{-1}$ to prioritize arm interactions kinematically. By training on $\mathcal{D}_{syn}$ with feasible trajectories synthesized from a single exemplar, the policy bypasses sim-to-real gaps while handling diverse bimanual coordination patterns.
Figure 2: The overview of BiDemoSyn. It consists of three stages (e.g., deconstruction, alignment, and optimization) based on a given demonstration. Then, we can apply our method to complete data collection efficiently and conveniently in real-world. It is best to zoom in to view the details.
4 Methodology
Here, we detail BiDemoSyn for synthesizing bimanual demonstrations $\mathcal{D}_{syn}$ from a single exemplar $\tau$ (teleoperated or kinesthetically guided) with three stages: Sec. 4.1 Deconstruction of One-Shot Teaching, which extracts invariant patterns and adaptable primitives; Sec. 4.2 Vision-Based Initial Frame Alignment, enabling generalization across geometric variations via efficient scene perception; and Sec. 4.3 Trajectory Modulation and Optimization, ensuring physical feasibility through hierarchical kinematic constraints. Each stage sequentially addresses scalability, adaptability, and real-world fidelity to bridge the gap between data efficiency and policy robustness. The overall framework is shown and illustrated in Fig. 2.
4.1 Deconstruction of One-Shot Teaching
Given a single demonstration $\tau\!=\!\{\textbf{o}_{t},\textbf{a}_{t}\}_{t=0}^{T}$, we decompose it into a bimanual execution blocks set $\{\mathcal{B}_{i}\}_{i=1}^{n}$, where each block $\mathcal{B}_{i}\!=\!(\textbf{s}_{i}^{L},\textbf{s}_{i}^{R})$ represents a discrete phase of dual-arm interaction. Here, $\textbf{s}_{i}^{L},\textbf{s}_{i}^{R}\!\in\!\mathbb{R}^{d}$ denote the left/right arm state sequences (including end-effector 6-DoF poses and gripper status) within block $\mathcal{B}_{i}$. Blocks are categorized based on motion coordination. For the single-arm motion, one arm exhibits significant state transitions while the other remains static. Formally, for threshold $\delta$ (minimal motion saliency) and $\zeta$ (static tolerance):
$\centering(\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\leq\!\zeta)\;\vee\;(\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\leq\!\zeta).\@add@centering$ (4)
While, for the dual-arm coordination, both arms synchronously or asynchronously adjust states:
$\centering\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\geq\!\delta.\@add@centering$ (5)
From Blocks to Atomic Execution Primitives (AEPs). To enable modular adaptation, each block $\mathcal{B}_{i}$ is further refined into a minimal motion unit AEP, where an arm undergoes a salient state transition (e.g., gripper closure, moving or pose shifts). For arm $A\!\in\!\{L,R\}$, an AEP is defined as:
$\centering\texttt{AEP}_{A}=\left\{\textbf{s}_{t}^{A}\rightarrow\textbf{s}_{t+\Delta t}^{A}\;|\;\left\|\textbf{s}_{t+\Delta t}^{A}-\textbf{s}_{t}^{A}\right\|_{2}\geq\gamma,\Delta t\leq T_{\textrm{max}},\texttt{CollisionFree}(\textbf{s}_{[t,\Delta t]}^{A})\right\},\@add@centering$ (6)
where $\gamma$ is a distance threshold, $T_{\textrm{max}}$ limits execution duration, and $\texttt{CollisionFree}(\cdot)$ enforces no collisions between the arm and objects/scene (optimized via subsequent forward kinematics). Unlike keyframe-based segmentation james2022coarse; shridhar2023perceiver; ma2024hierarchical, AEPs ignore acceleration profiles (assuming quasi-static motions) and prioritize task-oriented transitions over temporal granularity.
Semantic Categorization for Adaptive Synthesis. The final step categorizes refined blocks into invariant and variable types. The invariant blocks encode task-semantic primitives (e.g., screwing, pressing) or general motions (e.g., lifting, transferring), which remain structurally consistent across variations. The variable blocks, such as object-centric goal-conditioned grasping, adapt to instance-level geometric variations including object poses and shapes through:
$\centering\mathcal{B}^{\prime}_{i}=\mathcal{B}_{i}\circ\Phi(g,\textbf{o}_{\textrm{novel}}),\@add@centering$ (7)
where $\Phi(\cdot)$ is a vision-driven adapter, $g$ is the task goal, and $\textbf{o}_{\textrm{novel}}$ is the novel observation. This structured decomposition enables selective diversification, allowing variable blocks to adapt to new scenarios while preserving task fidelity in invariant blocks.
4.2 Vision-based Initial Frame Alignment
The variable blocks identified in Sec. 4.1 (e.g., initial grasping) require adapting to novel object poses and geometries. To achieve this, we propose a vision-driven adapter $\Phi(\cdot)$ that generalizes object-centric interactions in the given exemplar to new scenes. Given a novel observation $\textbf{o}_{\textrm{novel}}$, our method executes three steps: object perception, state estimation, and pose alignment, ensuring the initial robotic grasp aligns with the task goal $g$. Illustrations are shown in Fig. 3.
Figure 3: Illustrations of the initial frame alignment stage applied to tasks pouring (left and middle) and reorient (right). It shows that we can automatically adjust the grasp pose after the position, orientation and shape of the manipulated object changes.
Object Perception. We first detect and segment the target object in $\textbf{o}_{\textrm{novel}}$ using an open-vocabulary detector (e.g., YOLO-World cheng2024yolo or YOLOE wang2025yoloe) or vision foundation models (e.g., Florence2 xiao2024florence with SAM2 ravi2024sam for robustness to rare categories). In practice, we prioritize foundation models to avoid detection failures. Now let $M\!\subseteq\!\textbf{o}_{\textrm{novel}}$ denote the obtained object’s binary mask.
State Estimation. Instead of uitlizing category-level CAD-based 6D pose estimation models (e.g., FoundationPose wen2024foundationpose), we estimate the instance-level 6D object pose $\mathbf{P}_{\textrm{obj}}\!\in\!SE(3)$ using geometry-aware processing. Sepcifically, we adopt classic image moments chaumette2004image; kotoulas2007accurate to calculate the object mask centroid c and extract principal axes R:
$\centering\left\{\begin{array}[]{ccl}\textbf{c}&=&\frac{1}{|M|}\sum_{(u,v)\in M}(u,v,\textrm{d}(u,v))\\ \textbf{R}&=&\texttt{PCA}(\{(u,v,\textrm{d}(u,v))\;|\;(u,v)\in M\})\end{array}\right$.\@add@centering (8)
where $\textrm{d}(u,v)$ is the depth value. PCA is used to fit 3D points of $M$ to determine orientation $\textbf{R}\!\in\!SO(3)$. This yields $\mathbf{P}_{\textrm{obj}}\!=\!(\textbf{R},\textbf{c})$, robust to arbitrary object states (e.g., fallen, inverted).
Pose Alignment. Let $\mathbf{P}_{\textrm{demo}}$ denote the object pose in the given demonstration. We compute the rigid transformation $\mathbf{T}\!\in\!SE(3)$ that maps $\mathbf{P}_{\textrm{demo}}$ to $\mathbf{P}_{\textrm{obj}}$. This transformation is applied to the initial grasp pose $\mathbf{P}_{\textrm{grasp,demo}}$ in the variable block $\mathcal{B}_{i}$, yielding the adapted grasp pose:
$\centering\mathbf{P}_{\textrm{grasp,novel}}=\mathbf{T}\cdot\mathbf{P}_{\textrm{grasp,demo}}=(\mathbf{P}_{\textrm{obj}}\cdot\mathbf{P}^{-1}_{\textrm{demo}})\cdot\mathbf{P}_{\textrm{grasp,demo}}.\@add@centering$ (9)
The robot arm then executes a collision-free motion to reach $\mathbf{P}_{\textrm{grasp,novel}}$, ensuring goal-conditioned grasping without relying on task-irrelevant dense grasp proposals produced via 6-DoF grasp pose detectors fang2020graspnet; fang2023anygrasp. By integrating with the decomposition stage, the adapted grasp pose directly modifies the variable block $\mathcal{B}_{i}$ into $\mathcal{B}^{\prime}_{i}$, enabling synthesis of new trajectories.
4.3 Trajectory Modulation and Optimization
After deconstructing the demonstration into blocks (Sec. 4.1) and adapting variable blocks via vision-based alignment (Sec. 4.2), the synthesized trajectory requires two critical refinements to ensure executability and robustness as below:
Collision-Aware Validation. Each adapted block undergoes kinematic feasibility checks. First, we validate the reachability of target poses $\mathbf{P}_{\textrm{grasp,novel}}$ using Inverse Kinematics (IK), ensuring the robot joint limits are satisfied. If the result returned is singular or unreachable (rarely happens), we will delete this sample. Second, a motion planner chitta2012moveit; schulman2014motion verifies potential collision between arms and scene objects by solving for collision-free paths between the start and end states in each block. This simplified two-point constraint reduces computational overhead while preserving safety, assuming the trajectory of original given demonstration is collision-free.
Instance-Level Motion Adaptation. To handle geometric variations across object instances with differing shapes and sizes, we adjust motion primitives in variable blocks based on the object 3D bounding box dimensions. Let $\Delta l$, $\Delta w$ and $\Delta h$ denote the length, width and height differences between the novel object bounding box $\textbf{b}_{\textrm{novel}}$ and the original $\textbf{b}_{\textrm{demo}}$. Motion endpoints (e.g., grasp or release positions) are offset by:
$\centering\mathbf{s}_{i,e}^{A^{\prime}}=\mathbf{s}_{i,e}^{A}+\lambda(\Delta l,\Delta w,\Delta h),\@add@centering$ (10)
where $\lambda$ scales the adjustment (empirically set to $0.8\sim 1.0$), and $A\!\in\!\{L,R\}$. For irregular shapes, depth-aware masks $M$ extracted in Sec. 4.2 can approximate volumetric differences, or minimal human input provides precise measurements.
This hierarchical optimization stage operates at the block level rather than full trajectories, and enables efficient scaling while preserving the task semantics encoded in invariant blocks. Finally, validated and adapted blocks are recombined into synthetic trajectories $\tau^{\prime}\!=\!\{\tau_{\textrm{inv}},\{\mathcal{B}^{\prime}_{i}\}\}$, populating diverse and physically feasible demonstrations $\mathcal{D}_{syn}$.
5 Experiments
Our experiments address three core questions. $\mathsf{Q}1$: Is BiDemoSyn truly efficient and user-friendly? $\mathsf{Q}2$: Does BiDemoSyn enable scalable visuomotor imitation learning? $\mathsf{Q}3$: Does synthetic demonstrations generalize to spatial and object variations?
5.1 Experimental Setup and Protocol
Tasks. We evaluate on six bimanual manipulation tasks (refer Fig. 9) requiring contact-rich coordination: plugpen, inserting, unscrew, pouring, pressing and reorient. These tasks cover diverse primitive skills (e.g., grasping, rotating and handover), and involve both rigid and articulated objects. The one-shot demonstration for each task is collected via kinesthetic teaching. For the platform (refer Fig. 8), two fixed-base arms equipped with parallel grippers are used. Scene perception is provided by a stereo camera capturing binocular images. More details of all tasks and the corresponding data collection process are in Appendix Sec. A.
Policies & Baselines. For a comprehensive comparison, baselines contain two categories. One category is purely for data collection including point cloud editing (DemoGen xue2025demogen), real robot auto-rollout (YOTO zhou2025you), and human drag teaching (close to Teleoperation). Generally, the quality of collected demonstrations by these baselines is better in turn, but the cost is more time-consuming. After preparing the training data, we adapt three advanced visuomotor policies (DP chi2023diffusion, DP3 ze2024dp3 and EquiBot yang2024equibot) to bimanual settings by modifying their modeling spaces to dual-arm actions (refer Sec. 3). Observation inputs are RGB-only images or segmented 3D point clouds of task-relevant objects. Final trained policies operate in the open-loop discrete keyposes prediction to align with our synthesized demonstration format. The another category is directly for bimanual manipulation without retraining including the zero-shot ReKep huang2024rekep, an advanced ReKep+ with oracle-level grasp labels at the beginning, and one-shot ODIL wang2025one.
Metrics. We evaluate (1) synthesis efficiency: average time per demo over 50 tests, (2) data quality: visual authenticity of newly acquired observations, (3) success rates: real robot deployment with 30 trials per task, and (4) generalization: performance on unseen object placements and geometries.
Figure 4: (A) The data collection efficiency comparison of different baselines and our BiDemoSyn, and (B) the generated data quality illustration of DemoGen. Although DemoGen has the highest synthesis efficiency, it cannot avoid visual artifacts caused by perspective transformation, so its data quality is the lowest. Our method can achieve a balance between the speed and quality.
5.2 Comparison and Presentation of Results
We here answer three questions raised earlier, demonstrating that BiDemoSyn can efficiently synthesize high-quality real-world demonstrations, enabling scalable and generalizable visuomotor policy training with minimal human input. Then there are some snapshots of real robot execution effects.
($\mathsf{A}1$): The efficiency and usability of BiDemoSyn have obvious advantages over baselines. Without bells and whistles, our BiDemoSyn demonstrates superior efficiency and usability as illustrated in Fig. 4. It requires about 5 seconds per demonstration to synthesize new trajectories, where primarily involving manual object repositioning and visual verification of initial frame alignment, followed by the trajectory optimization. In contrast, DemoGen generates trajectories in less then 1 second via scripted edits but suffers from perspective distortion artifacts, degrading data quality. The specific qualitative effect can be compared with Fig. 4A and our results in Fig. 2 and Fig. 9. By relying on manual alignment and auto-replay, YOTO takes 42 seconds per demo which is equivalent to real robot execution time. While, teleoperation demands 91 seconds per demo for near-perfect but labor-intensive collection. These results highlight BiDemoSyn’s unique balance of speed (automated optimization) and reliability (artifact-free synthesis), making it the only method scalable to thousands of demonstrations without compromising real-world fidelity.
Table 1: Quantitative comparison results of success rates for different methods under in-distribution and out-of-distribution evaluations. The # means the BiDemoSyn training data size.
  in-distribution evaluations out-of-distribution evaluations
Policy
(Input
Mode)
Method
plugpen
inserting
unscrew
pouring
pressing
reorient
Average
Success
Rate
plugpen
inserting
unscrew
pouring
pressing
reorient
Average
Success
Rate
#3888 #7776 #1152 #2592 #1296 #1008 #2916 #3888 #1008 #972 #324 #756
Training
-Free
ReKep 15/30 13/30 12/30 15/30 12/30 09/30 42.2% 12/30 10/30 10/30 09/30 09/30 06/30 31.1%
ReKep+ 17/30 17/30 14/30 19/30 20/30 11/30 54.4% 15/30 14/30 12/30 12/30 11/30 09/30 40.6%
ODIL 18/30 19/30 13/30 18/30 18/30 12/30 54.4% 12/30 14/30 08/30 11/30 13/30 08/30 36.7%
DP
DemoGen 18/30 20/30 16/30 16/30 14/30 15/30 55.0% 08/30 03/30 10/30 04/30 03/30 07/30 19.4%
YOTO 19/30 20/30 16/30 17/30 14/30 16/30 56.7% 08/30 04/30 10/30 05/30 04/30 07/30 21.1%
BiDemoSyn 22/30 24/30 22/30 21/30 17/30 19/30 67.8% 13/30 10/30 18/30 15/30 08/30 12/30 42.2%
DP3
DemoGen 21/30 24/30 19/30 20/30 18/30 17/30 66.1% 11/30 06/30 14/30 07/30 05/30 11/30 30.0%
YOTO 22/30 23/30 20/30 20/30 18/30 19/30 67.8% 12/30 06/30 15/30 08/30 06/30 14/30 33.9%
BiDemoSyn 26/30 28/30 25/30 24/30 21/30 22/30 81.1% 16/30 14/30 21/30 18/30 11/30 18/30 54.4%
EquiBot
DemoGen 22/30 24/30 20/30 20/30 19/30 19/30 68.9% 13/30 10/30 16/30 11/30 10/30 13/30 40.6%
YOTO 23/30 24/30 20/30 21/30 20/30 19/30 71.1% 15/30 11/30 18/30 14/30 12/30 16/30 47.8%
BiDemoSyn 28/30 28/30 26/30 25/30 24/30 25/30 86.7% 18/30 20/30 24/30 21/30 17/30 20/30 66.7%
Figure 5: Comparison between training data scale and the success rate. Less sized data is randomly sampled out of the total dataset at the task level. DP3 is chosen as the visuomotor policy.
($\mathsf{A}2$): Demonstrations obtained via BiDemoSyn can support scalable imitation learning. We applied BiDemoSyn for efficient collection of thousands demonstrations per task (Fig. 9 right), densely covering the workspace with instance-level object diversity, thereby providing abundant training data for imitation learning. For two reproduced baselines, DemoGen xue2025demogen matches our data scale but suffers from visual artifacts, while YOTO zhou2025you collects only 1/10 the data per task for being limited by its time-consuming replay mechanism. After obtaining adequate data, we trained three advanced visuomotor policies DP chi2023diffusion, DP3 ze2024dp3 and EquiBot yang2024equibot. As shown in Tab. 1 left, BiDemoSyn achieves superior in-distribution (ID) success rates across all six tasks, outperforming both baselines regardless of the policy architecture. This confirms that BiDemoSyn’s artifact-free synthesis and efficient data scaling reliably support the scaling laws of imitation learning. Besides, it is unsurprising that policies trained with demonstrations consistently outperform the two training-free baselines ReKephuang2024rekep and ODIL wang2025one. To further quantify scalability, we evaluated training data size-performance relationships on all six tasks of all methods. As shown in Fig. 5, BiDemoSyn-trained policies exhibit stronger scaling trends than DemoGen (e.g., 60% vs. 39% at 1106 demos), aligning with community observations that data quality and quantity jointly drive visuomotor policy performance liu2025rdt; black2024pi0; pertsch2025fast; lin2025data.
Figure 6: Analysis of spatial generalization, including variations in (A) position sampling density and (B) orientation sampling diversity. DP3 is chosen as the visuomotor policy.
($\mathsf{A}3$): Policies trained on BiDemoSyn data can achieve better generalization to unseen variations. To evaluate the generalization to novel instances, we conduct controlled tests across all six tasks. For each task, we exclude demonstrations associated with one or a pair of randomly selected objects (e.g., an arbitrary bottle in unscrew or a bottle-cup pair in pouring) from the training set, resulting in training sizes of 2916/3888/1008/972/324/756 demos for six tasks, respectively. These excluded objects are reintroduced exclusively during real-world testing, with identical data splits applied to DemoGen and YOTO baselines for fair comparison. As shown in Tab. 1 right, policies trained on BiDemoSyn data consistently achieve higher success rates compared to all baselines in out-of-distribution (OOD) settings. While all methods exhibit performance drops relative to ID evaluations, BiDemoSyn’s superior OOD robustness highlights its ability to capture task-invariant features through vision-aligned synthesis. To further analyze spatial generalization, we vary the density of positional and orientation coverage in training data: for unscrew, we synthesize trajectories with sparse-to-dense workspace coverage (Fig. 6A), and for reorient, we incrementally increase the diversity of object orientations (Fig. 6B). To exclude the effect of data size differences, the total number of train-set is always augmented to a standard size (e.g., #1152 and #1008) via the DemoGen to edit some observations at small distances. Finally, policies trained on BiDemoSyn data with dense spatial coverage (e.g., $10\times 10$ workspace coverage in unscrew) achieve 76.7% success on unseen positions, compared to 53.3% for sparse coverage ($6\times 6$ grid cells). Similarly, orientation diversity in reorient improves generalization from 37.3% (lower diversity with #Ornt.=3) to 73.3% (higher diversity with #Ornt.=7). These results confirm that BiDemoSyn’s ability to cheaply synthesize positionally and orientationally diverse demos directly translates to enhanced policy generalization, aligning with empirical insights that broad data coverage mitigates spatial distribution shifts.
Figure 7: Visualization of all six bimanual tasks performed on real robots. All models are trained and tested under the ID evaluations. EquiBot is chosen as the visuomotor policy. Key dual-arm coordination movements associated with each task are partially enlarged for quick review.
Qualitative Results: Real rollouts in Fig. 7 showcase effectiveness across six tasks. For example, the plugpen policy precisely aligns and inserts the pen into the narrow cap even with $\pm$5mm positional noise. In inserting, dual-arm coordination achieves rough but proximate peg-in-hole precision. The unscrew demonstrates robust rotational synchronization, successfully loosening lids with variable thread tightness. The pouring policy adapts liquid flow control to container geometries, minimizing spills despite tilted orientations. For pressing, predicted actions reliably activate nozzles without over-pressuring, while reorient handles object handovers and flips with smooth dual-arm transitions. These visualizations corroborate our quantitative findings, emphasizing BiDemoSyn’s ability to handle real-world complexity in contact-rich manipulation.
6 Conclusion and Limitation
We present BiDemoSyn, a framework for synthesizing real-world bimanual demonstrations from a single human example. By decomposing tasks into invariant and adaptable components, leveraging vision-guided scene alignment, and hierarchically optimizing trajectories, BiDemoSyn generates diverse, collision-free data without synthetic gaps. Experiments on six complex contact-rich tasks demonstrate its efficiency, scalability, and generalization, outperforming strong baselines and pure teleoperation. This work establishes a new paradigm for scalable imitation learning, bridging the divide between data quantity and physical fidelity in bimanual robotic manipulation.
Limitations: While BiDemoSyn excels in static, geometrically varied scenes, it faces challenges in dynamic environments and extreme shape variations (e.g., deformable fabric objects). The current hierarchical optimization assumes quasi-static motions, limiting its applicability to highly dynamic tasks like catching. Additionally, synthesizing trajectories for multi-stage tasks with interdependent contacts (e.g., lacing ropes) requires further extension. Future work will integrate dynamic perception and adaptive contact modeling to address these constraints.
Appendix
This supplementary part provides detailed explanations and extensions to the main paper. Sec. A outlines the design rationale for our six bimanual manipulation tasks, including platform specifications, task-specific objectives, and one-shot demonstration acquisition protocols. Sec. B details the data collection standards, UI-assisted collection tools, and post-processing workflows. These constitute the trajectory synthesis pipelines for generating real-world demonstrations with BiDemoSyn. We have also added more details about BiDemoSyn to help readers better understand this study. Sec. C elaborates on policy training and deployment, covering implementation specifics for baseline methods (DemoGen xue2025demogen, YOTO zhou2025you) and diffusion policies (DP3 ze2024dp3, EquiBot yang2024equibot), alongside real-robot deployment. Sec. D provides additional visualizations of real-robot executions and failure case analyses, offering insights into current limitations. Sec. E concludes with reflections and future directions, aiming to extend BiDemoSyn to challenge more complex tasks (e.g., deformable or dynamic manipulation) and enhance cross-domain generalization. Sec. F is the statement of the use of Large Language Models.
Appendix A Design of Bimanual Manipulation Tasks
Figure 8: Left: The fixed-base dual-arm manipulator platform (a table with two robot arms, two grippers and the binocular camera) used in this research. Right: The object assets involved in our six bimanual manipulation tasks. All objects have been scaled down proportionally.
A.1 Hardware and Platform
[…]
A.2 Bimanual Task Formulation
We design six bimanual manipulation tasks (including plugpen, inserting, unscrew, pouring, pressing and reorient) to comprehensively evaluate dual-arm coordination and generalization. Each task involves at least two category-level object instances (refer Fig. 8 right) and requires diverse primitive skills, such as single-arm actions (grasping, placing, carrying, precision rotation) and dual-arm coordination (fixed-twisting, plug-in, handover, alignment). All tasks are long-horizon, starting from goal-conditioned initial grasping (objects fully separated from robots), contrasting prior works already grasping/holding the manipulated object and focusing on short-horizon atomic skills such as untwisting lin2024twisting; bahety2024screwmimic or handover huang2023dynamic. Fig. 9 shows some examples. Below we detail each task:
[…]
  • plugpen: plug the pen with the pen cap. States: A marker pen body and its cap are placed separately on the table. The pen body lies horizontally with its tip roughly pointing to the right arm, while the cap is also positioned horizontally with its socket facing the left arm. Steps: Left arm grasps pen body; right arm grasps cap. Arms lift and spatially align the pen tip with the cap’s socket before insertion. Challenges: Directional grasping alignment, sub-millimeter positional tolerance for insertion.
  • inserting: insert the marker pen into the cup. States: An empty handle-free cup is placed upside-down on the table, and a closed marker pen lies horizontally nearby. Steps: Left arm flips the cup (about 180 rotation); right arm reorients the marker vertically. Arms coordinate to align and insert the marker into the upright cup. Challenges: Dual-arm rotational synchronization, tight insertion tolerance ($\pm$2mm).
  • unscrew: open the bottle by twisting cap counterclockwise. States: A translucent plastic bottle (filled with colorless water) stands upright on the table, with its cap tightly screwed on. Steps: Left arm stabilizes the bottle; right arm grasps the cap and rotates it vertically (with multiple degree-fixed rotations). Challenges: Force-agnostic cap grasping (no torque sensing), rotational precision ($\pm$5 per turn).
  • pouring: pour water from the bottle into the mug cup. States: An uncapped water bottle (about $3/4$ full) and an empty handled mug are placed on opposite sides of the table. Steps: Left arm tilts the bottle (about 90); right arm positions the mug to catch water. Challenges: Fluid dynamics approximation, nozzle-cup alignment under gripper deflection.
  • pressing: press bottle and catch water using the cup. States: A shampoo bottle (with a pressable nozzle) containing water and an upright empty handle-free cup are placed apart. Steps: Right arm vertically presses the nozzle; left arm tilts the cup to catch water. Challenges: Nozzle positioning accuracy ($\pm$5mm), open-loop force control.
  • reorient: flip the spoon so that the bottom is facing up. States: A long spoon or shovel (metal or plastic) lies concave-down on the table, with randomized orientation and position. Steps: Right arm grasps the spoon center, reorients it mid-air, and hands it to the left arm, which then flips and places it. Challenges: Unstable grasp during handover, rotational precision for flipping ($\pm$10).
These tasks collectively stress spatial reasoning, contact-aware coordination, and generalization to instance-level variations, which are cornerstones of real-world bimanual manipulation. Besdies, it is important to note that, we have shared hardware constraints: (1) No force/torque sensing limits contact-rich adjustments (e.g., pen plug-in force, cap twisting force, or pressing force). (2) Binary gripper states (open/closed) restrict fine-grained manipulation (e.g., plastic bottles are always deformed by excessive clamping). (3) Third-person view occlusions occasionally hinder precise alignment (e.g., cannot discern the groove side of the pen cap).
A.3 Obtaining and Decomposing the One-Shot Demonstration
As stated in the main content, we collect one-shot demonstrations via kinesthetic teaching: an operator manually guides both arms through task-critical waypoints, recording the 6DoF end-effector poses (relative to each robot base frame) and gripper binary states at each pause. Objects are placed in fixed initial configurations (allowing small positional tolerance) to ensure consistency. The recorded waypoints are then executed autonomously by the robot control API, which solves inverse kinematics between consecutive given poses and synchronizes gripper actions (e.g., closing after reaching a pre-grasp pose). During auto-execution, stereo camera observations (10Hz) and dual-arm joint/end-effector states are logged. For the real rollout effect of the one-shot demonstration related to each task, please refer to our Supplementary Videos. Finally, demonstrations are deconstructed into task-aware blocks to support subsequent trajectory synthesis.
To enable reliable downstream synthesis, the one-shot kinesthetic demonstration must be collected in a way that naturally exposes task-relevant structure. During kinesthetic teaching, the demonstrator physically guides the two manipulators through the task while pausing at intuitive semantic boundaries, such as after a stable grasp, after completing an alignment adjustment, or before switching from transportation to interaction. These pauses, which arise organically from human motion patterns, produce clear temporal discontinuities that align well with the block-based representation introduced in Sec. 4.1 of the main text. Importantly, this process does not require any manual annotation or labeling. The demonstration is captured in a continuous stream exactly as in standard kinesthetic teleoperation, and the boundaries between coarse blocks emerge directly from the dynamics of human-guided motion. This makes the procedure no more labor-intensive than collecting a single ordinary demonstration.
We explored alternative strategies for fully automatic segmentation, such as keyframe/keypose detection and curvature-based change-point identification james2022coarse; shridhar2023perceiver; ma2024hierarchical. However, reliably identifying semantically meaningful waypoints from a single demonstration proved brittle, especially when waypoints must correspond to subtle but crucial phases of contact establishment, re-grasping, or dual-arm synchronization. Hardware-assisted approaches (e.g., instrumented gloves as in DexCap wang2024dexcap or end-effector teaching interfaces as in UMI chi2024universal and DexUMI xu2025dexumi) can offer additional cues for segmentation but introduce alignment overhead, cross-morphology calibration challenges, and potential loss of semantic correspondence between human motion and robot embodiment. Given these trade-offs, the kinesthetic-teaching-driven segmentation remains a pragmatic and robust choice: it imposes no additional annotation burden, preserves semantic alignment across blocks, and provides boundaries well suited for constructing invariant and variant components for real-world synthesis.
Appendix B Details of Data Collection, Processing and Synthesis
Figure 10: The specific grid cell division way for each task. For tasks involving two manipulated objects, the total number of grid cells will be divided equally between the left side and right side.
Figure 11: UI interface examples. When collecting data, the six tasks need to display different grid cells in real time, as well as perceive and track objects related to the tasks. Note that these points, crosses and lines are drawn digitally, which are not marks in the real world.
B.1 Task-Oriented Collection Standards
[…]
B.2 Collection UI Interface Development
To streamline data collection, we develop a real-time UI interface (Fig. 11) that visualizes task-specific grid cells, dynamically tracks object positions/poses, and validates alignment with the one-shot demonstration. The interface overlays grid layouts onto live camera feeds, highlights detected objects (using YOLOE wang2025yoloe for fast detection and segmentation), and marks completed cells using red crosses. The position of objects is represented by red dots (some objects also use yellow arrows to indicate their orientation). In addition, for cups and bottles that do not need to pay attention to changes in orientation, we did not use the mask centroid as their representive point during collection, but instead used the lowest endpoint in contact with the desktop. A sidebar displays positional/orientational deviations from the reference demonstration, enabling quick verification. If object perception fails, the system falls back to Florence2 xiao2024florence and SAM2 ravi2024sam for robust detection and segmentation. Operators trigger a ”capture-align-validate-save” cycle every 2 seconds. And successful saves prompt visual feedback (UI flash of corresponding grid cells), signaling readiness for the next sample. This setup eliminates manual grid marking and minimizes human idle time, achieving a high speedup while ensuring spatial consistency. Please refer our Supplemental Videos that demonstrate the vivid workflow, illustrating how the UI balances efficiency with precision.
B.3 Post-Processing of Collected Data
This part includes the last two stages of BiDemoSyn: initial frame alignment and trajectory optimization (defined in the method part of main paper). First, pre-recorded positional/orientational deviations (validated via the UI interface during collection) align novel observations with the one-shot demonstration reference frame using the rigid transformation. Next, trajectory optimization ensures kinematic feasibility. We prune singular configurations and validate collision-free paths using the robot IK solvers and motion planners. Over 96% of trajectories pass validation, with failures (e.g., unreachable corner grid cells) either re-collected or discarded. Finally, validated trajectories are stored in $\mathcal{D}_{syn}$ to complete the synthesis pipeline.
B.4 Clarification on Hyperparameters and Predefined Symbols
In Sec. 4, several task-specific thresholds and predefined parameters are introduced for clarity and reproducibility. Here we further explain their meaning and selection strategy. For instance, threshold symbols $\delta$, $\zeta$ and $\gamma$ in Eqn. 4, Eqn. 5 and Eqn. 6 are predefined only for one-shot deconstruction. The symbol $\delta$ is a minimal motion saliency threshold, used to identify boundaries between coarse blocks. The symbol $\zeta$ is a static tolerance threshold to segment motion boundaries under quasi-static assumptions. The symbol $\gamma$ is a refinement threshold used to create fine-grained AEPs after bounding block duration ($T_{\textrm{max}}$). And the scaling factor $\lambda$ in Eqn. 10 controls the adaptation from the canonical one-shot demonstration to new object instances with varying sizes. Theoretically, $\lambda\in\mathbb{R}^{+}$, yet empirical tuning shows that values within $[0.8,1.0]$ are sufficient for stable grasp transfer without destabilizing contact. Larger or smaller values tend to shift grasp points away from functional regions (e.g., bottle mid-section in unscrew), while $\lambda=1.0$ already suffices for most tasks. Similarly, thresholds for collision-checking or pose-alignment rejection (e.g., rejecting infeasible IK solutions or near-singular configurations) were determined through small-scale grid search and validated in pilot runs. Importantly, such hyperparameters are not per-trial tunable knobs, but rather fixed system constants, ensuring consistent data synthesis across tasks.
B.5 Additional Evaluation of Intermediate Processes
We provide further evaluation to justify intermediate design choices from three aspects: Segmentation Quality, Pose Estimation Robustness and Impact on Data Reliability. (1) Firstly, object masks were obtained from Florence2 xiao2024florence and SAM2 ravi2024sam, two state-of-the-art vision foundation models (VFMs). On held-out captures without foreground occlusion, the segmentation accuracy exceeded 97%, ensuring reliable geometric inputs for subsequent processing. (2) Secondly, the axis alignment for objects with a distinct orientation was based on classical image-moment theory: first-order moments define centroids, second-order moments define orientation. This method, given reliable masks, yields highly stable results. Quantitative inspection of $\sim$200 sampled trials revealed <<2% pose estimation errors, and only under severe occlusion. (3) Thirdly, in Sec. D.2 (Statistics and Analysis of Failed Cases), we showed that perception and orientation errors together account for  47% of failures, highlighting their importance. Nevertheless, these intermediate modules are already close to optimal under current open-loop constraints, which explains the high success rate ( 98%) of the data collection system.
B.6 Positioning with Respect to Prior Work
To avoid potential confusion with prior works such as ReKep huang2024rekep and ODIL wang2025one, we emphasize both the similarities and the fundamental differences. Like these methods, BiDemoSyn exploits one-shot demonstrations and compositional reasoning; however, its core objective is not direct policy execution, but rather rapid and scalable synthesis of real-world training demonstrations. This distinction is crucial: while prior methods target zero-/few-shot policy generalization (often trading reliability for flexibility), BiDemoSyn deliberately focuses on maximizing data quality and efficiency, enabling downstream training of robust visuomotor policies (e.g., DP3 ze2024dp3 and EquiBot yang2024equibot). Our evaluation against DemoGen xue2025demogen and YOTO zhou2025you, which are two strong baselines addressing the same bottleneck, demonstrates that BiDemoSyn achieves a superior balance between efficiency and data reliability, which is the practical need for scaling imitation learning.
A concurrent work MoMaGen li2025momagen tackles scalable bimanual mobile manipulation by generating diverse trajectories in simulation, focusing on base reachability and camera visibility through constrained optimization. Although sharing a high-level goal of reducing human effort, MoMaGen and BiDemoSyn operate under fundamentally different assumptions: MoMaGen relies on noise-free simulation assets and virtual scene perturbations, whereas BiDemoSyn performs real-world demonstration synthesis, directly addressing physical-domain challenges such as noisy perception, contact feasibility, and one-shot decomposition under real sensor observations. As a result, MoMaGen expands diversity through simulated scene variation, while BiDemoSyn enables physically grounded generalization across spatial and category-level object variations without sim-to-real transfer. Thus, BiDemoSyn should be viewed not as a direct competitor to simulation-based frameworks like MoMaGen, but as a complementary real-world solution specifically designed for scalable, high-fidelity data generation where simulation pipelines cannot fully capture real-world physical constraints.
B.7 Evaluation of Demonstration Quality
Evaluating the quality of synthesized demonstrations is essential because the downstream visuomotor policies depend on both the fidelity of observations and the physical feasibility of actions. We assess demonstration quality along two orthogonal dimensions that align with the structure of visuomotor imitation learning:
Visual Fidelity. This dimension evaluates whether the observations included in a synthesized demonstration faithfully reflect the real physical scene without introducing artifacts. Teleoperation and BiDemoSyn preserve raw RGB-D observations, ensuring perfect consistency with the physical environment. YOTO zhou2025you and DemoGen xue2025demogen modify the seed point cloud to reposition objects, which introduces geometric distortions and depth inconsistencies, particularly as the displacement grows (see Fig. 4B). These artifacts degrade the policy’s perception module by coupling unrealistic geometry with real sensor noise.
Trajectory Feasibility and Physical Grounding. A demonstration is considered high-quality if the associated action sequence: (1) yields a collision-free, dynamically stable trajectory, (2) respects the contact geometry implied by the task semantics, and (3) can be executed on the real robot without additional correction or replanning. Teleoperation and YOTO zhou2025you provide ground-truth trajectories. BiDemoSyn produces partially reused (invariant blocks) and partially adapted (variant blocks) trajectories aligned with real-world object pose and geometry. DemoGen xue2025demogen synthesizes trajectories directly from edited scene representations, which can deviate from physically stable contact geometries.
Finally, Fig. 4A reflects the interplay of these two components. Demonstrations with high-quality RGB-D observations and physically feasible actions rank highest. This ordering also predicts downstream policy performance: policies trained with cleaner, physically grounded demonstrations demonstrate higher success rates and better robustness under OOD evaluation. These expanded analyses clarify the qualitative and quantitative factors driving the reported comparisons and support the role of demonstration quality as a key determinant of real-world imitation learning performance.
B.8 Diversity of Synthetic Demonstrations
A natural question is whether additional diversity (such as reordering or randomizing block sequences) would improve the effectiveness of the synthesized dataset. In BiDemoSyn, we intentionally do not introduce such sequence-level variations. Our empirical findings align with recent analyses in large-scale teleoperation datasets focusing on the diversity exploration and ablation for scalable robotic manipulation shi2025diversity, which show that low-level behavioral variability (e.g., minor differences in grasp location, end-effector speed profiles, or redundant micro-motions) does not necessarily benefit policy learning and may even hinder convergence. Instead, we adopt the principle that task success is primarily determined by achieving the correct functional interactions, not by how those interactions are executed.
Accordingly, in the design of BiDemoSyn, we maintain the original causal structure of the task by preserving the block ordering extracted from the one-shot demonstration, while introducing diversity along the dimensions that matter for generalization, namely: (1) spatial variation of object pose (translation + rotation), and (2) category-level instance variation (shape and size differences within the same object class). These forms of variation are directly tied to visuomotor generalization and produce clear benefits in downstream policy performance (see Tab. 1). In contrast, randomizing the internal block structure introduces behavioral diversity that is orthogonal to task semantics and does not improve success rates in our preliminary tests. We therefore avoid unnecessary variation and instead focus on environment-induced diversity, which is both meaningful and consistent with the underlying task geometry.
Appendix C Training and Deployment of Imitation Policies
C.1 Reproduction of Baselines
[…]
Figure 12: The adapted visuomotor policy network architecture for imitation learning of bimanual manipulation tasks. The input visual observation is simplified to a point cloud of manipulated objects. Note that the vision encoder is slightly different in DP3 ze2024dp3 and EquiBot yang2024equibot. The illustrated one is for EquiBot.
C.2 Implementation of Imitation Policies
Both DP3 ze2024dp3 and EquiBot yang2024equibot (refer Fig. 12) ingest segmented 3D object point clouds but differ in network architectures. Specifically, DP3 employs specially designed multiple MLPs to encode point clouds into latent features, then directly regressing dual-arm keyposes. EquiBot enhances robustness via a SIM(3)-equivariant PointNet variant: point clouds undergo rotation/scale-invariant feature extraction before action prediction. This equips EquiBot with improved generalization to unseen object variations (refer Tab. 1 in the main paper), achieving 5–12% higher success rates than DP3 under identical training data and settings. Both policies operate in open-loop, aligning with synthesized demonstration formats. Below, we give a more detailed description of the visuomotor policy modification for the dual-arm action prediction.
  • Spaces of observation and action. We adopt a 13-dimensional proprioception vector and a 7-dimensional action space for each robot arm, respectively. The proprioception data for each arm consists of the following information: a 3-dimensional end-effector position, a 6-dimensional vector denoting end-effector orientation (represented by two columns of the end-effector rotation matrix), a 3-dimensional vector indicating the direction of gravity, and a scalar that represents the degree to which the gripper is opened. The action space for each arm consists of the following information: a 3-dimensional vector for the end-effector position offset, a 3-dimensional vector for the end-effector angular velocity in axis-angle format, and a scalar denoting the gripper action. For all bimanual tasks, the observation horizon is set to 1, so we only use the initial state observation of left arm as one of the network inputs. And the initial state of right arm is always fixed in each task. For the number of action steps (also the length of the predicted horizon), we simplify it and set the prediction length to the number of keyposes $K$, which can be extracted from the start and end actions of each Atomic Execution Primitives (AEPs) during task deconstruction.
  • Network architecture. For DP3, it is a variant of Diffusion Policy chi2023diffusion with a simpler point cloud encoder. It also designs a two-layer MLP to encode robot proprioceptive states before concatenating with the observation representation. For EquiBot, we use a SIM(3)-equivariant PointNet++ yang2024equivact with 4 layers and hidden dimensionality 128 as the feature encoder. For the noise prediction network, we inherits hyperparameters from the original Diffusion Policy. Specifically, to optimize for inference speed in all experiments, we use the DDIM scheduler song2021denoising with 8 denoising steps, instead of the DDPM scheduler ho2020denoising which performs up to 100 denoising steps.
  • Sampling of point cloud. As we all known, setting the number of points to sample in the point cloud observation is a key hyperparameter to consider when designing an architecture that takes point cloud inputs. In our experiments, we found out that using 1024 points is sufficient for all tasks and policies. In particular, we have tried increasing the number of point clouds to 2048 or more, but the evaluation improvement in each task is minimal, and this will also cause the storage occupied by the training observation data to be too large and the training time cost to increase. Therefore, reducing the number of points to 1024 can make training faster without hurting performance. And all our policy models can be trained on a GeForce RTX 3090 Ti with 24 GB of memory.
  • Training and evaluation. When conducting the in-distribution (ID) and out-of-distribution (OOD) evaluations with full demonstrations, we train all methods for 500 and 1,000 epochs, respectively. Otherwise, when using the 1/10 training data (especially the YOTO), we train all methods of the ID and OOD evaluations for 2,000 and 4,000 epochs, respectively. For all experiments, the batch-size is always set to 64. We only evaluate the last one checkpoint saved at the end of training. For every evaluation in the real world, we run the policy in a randomly initialized placement of objects, and record the average success rate achieved by the policy.
Figure 13: Qualitative real rollout samples from the left RGB camera for all evaluation scenarios. From top to bottom, they are defined six bimanual tasks. Best to view after zooming in.
In addition, we have adopted the object-centric point cloud input. At inference time, we also need to preprocess the binocular RGB observations to obtain the point cloud of manipulated objects. This core design relies on the still rapidly developing capabilities of vision foundation models (VFMs). Here we leverage SOTA open vocabulary detection method Florence-2 xiao2024florence and segmentation method SAM2 ravi2024sam to automatically extract object masks and then filter out corresponding point clouds. Despite this, occasionally we may fail to segment desired objects accurately, and in these special cases we will manually correct the masks. These cases are not counted as failed evaluation trails due to not involving significant elements of bimanual robot manipulation. Because we believe that the next generation of VFMs can alleviate these problems, or we can directly address them through domain adaptation, test-time adaptation, or adjusting input prompts.
Appendix D More Real Robot Results Visualization and Analysis
D.1 More Examples from Real Rollouts
As shown in Fig. 13, it supplements more qualitative results with additional real-robot execution visualizations, highlighting nuanced state transitions and task-specific challenges. These visualizations underscore BiDemoSyn’s ability to handle real-world complexities (such as mechanical tolerances and imperfect perception), while also exposing limitations in dynamic force modulation (e.g., over-pressing bottles or lids). Please refer our Supplemental Videos which provide frame-by-frame analyses of these executions, further dissecting critical phase and offering insights into policy decision-making under uncertainty.
D.2 Statistics and Analysis of Failed Cases
While policies trained with BiDemoSyn achieve high success rates, failure analysis reveals systematic challenges. Note that although the policy we trained is executed end-to-end (it establishes an implicit mapping from observations to action outputs), we can still align the analysis from the stage where its failure cases are located to find the core cause of the error. Finally, using the experimental results of EquiBot under out-of-distribution evaluations, we categorize failures into five types according to the task execution logic (mainly from the design module of BiDemoSyn):
Figure 14: Typical failure cases. Left four cases: These include two cases from the plugpen task that incorrectly detected the marker body, and two cases from the reorient+unscrew task that incorrectly estimated the lying bottle’s orientation. Middle six cases: These include three cases from the pouring task where water was spilled or the mug cup was misgrabbed, and three cases from the unscrew task where the gripper failed to grab the bottle’s body or cap accurately (these are extra footages from a third-person perspective). Right four cases: These include two cases in the inserting task where the robotic arm’s end effector occasionally obscures the marker pen, and two cases in the reorient task where human interference causes the hand to partially obscure the spoon/shovel. Here, visual perceptual defects do not necessarily lead to task failure, demonstrating the robustness of the proposed orientation estimation algorithm.
  • Visual Perception Errors (15%): Inaccurate object segmentation (e.g., mistaking the tabletop as part of the marker pen) leads to misaligned grasps.
  • Orientation Estimation Errors (32%): Incorrect object pose estimation (e.g., misjudging a spoon’s concave direction) causes malformed trajectories (e.g., flipping to the wrong side).
  • Alignment Failures (18%): Vision-guided initial frame misalignment (e.g., bottle nozzle offset by >>1cm) propagates errors to subsequent steps (e.g., spilling during pouring).
  • Initial Grasp Failures (28%): Collisions during pre-grasp motions (e.g., gripper nudging objects) or unstable grasps (e.g., slippage in reorient) prevent task initiation.
  • Error Propagation (7%): Cumulative errors from earlier stages compound into irreversible failures (e.g., slightly misaligned pluging between pen body and pen cap in plugpen).
As can be seen, the orientation estimation and initial grasp failures dominate, reflecting two core challenges: (1) current pose estimators struggle with symmetric or textureless objects (e.g., metal spoon or shovel), and (2) gripper-centric path planning lacks fine-grained contact modeling (e.g., avoiding pre-touch collisions for irregular shapes). Addressing these requires advances in category-agnostic pose estimation and short-horizon contact optimization, which are critical directions for our future work. Some visual failure examples are shown in Fig. 14, which provides a foundation for a full understanding of the BiDemoSyn system and its robust adaptation to allow for dynamic pre-grasping. More dynamic rollouts can be found in our Supplementary Videos.
D.3 Two More Complex, Long-Horizon Tasks
To further evaluate the scalability of BiDemoSyn beyond the six primary bimanual tasks, we introduce two additional long-horizon tasks that require sequential manipulation steps and accumulate significantly more physical error throughout the execution:
  • reorient+unscrew: The robot must first upright a lying down bottle (reorient phase), then perform a dual-arm unscrewing motion to remove the cap (unscrew phase). This task stresses both precise estimation of bottle orientation and robust reuse of multi-step, contact-rich interaction patterns.
  • unscrew+pouring: The fisrt phaset is the original unscrew task. After unscrewing the bottle cap, the robot must lift a mug cup, perform dual-arm coordination to tilt the bottle, and pour water into the cup. This composition integrates two high-precision subtasks and requires stable arm coordination across multiple transitions.
Figure 15: Qualitative real rollout samples from the left RGB camera for two newly added complex long-horizon bimanual tasks. Best to view after zooming in.
Data Synthesis Using BiDemoSyn. These two newly added tasks were synthesized following the same three-stage pipeline described in the main text. We briefly summarize the configurations: for reorient+unscrew, we collected initial observations for 4 bottle instances, each placed at 7 orientations across 36 grid cells, yielding 1008 synthesized demonstrations; for unscrew+pouring, we collected initial observations for 2 bottle and 2 cup instances, each placed on 18 grid cells, yielding 1296 synthesized demonstrations. Importantly, both tasks introduce longer temporal horizons, sequential coordination, and increased compounding error, thus providing a stronger stress test than the original six tasks.
Policy Training and Performance. Using the synthesized datasets, we trained visuomotor policies for each long-horizon task based on the adapted EquiBot architecture (the strongest performer in Tab. 1 of the main text). Across both tasks, the trained policies exhibited high robustness to spatial variation (different placements and bottle orientations), reliable execution of multi-step dual-arm coordination, and consistent task success under mild external perturbations introduced during pre-grasp interpolation. Fig. 15 presents some qualitative results with additional real-robot execution visualizations of these two new tasks. And Fig. 16 provides illustrative examples of continuous dynamic interference scenarios among tasks pouring and reorient+unscrew. Full dynamic demonstrations and rollouts of all eight tasks (previous 6 tasks and the new 2 tasks) are included in the Supplemental Videos. These results further validate BiDemoSyn’s ability to rapidly scale to new, long-horizon tasks with minimal additional engineering, and reinforce its suitability as a practical framework for synthesizing large-scale, real-world training data.
Figure 16: Qualitative examples of applying dynamic interferences during the pre-grasping execution. From top to bottom, they are segments of the dynamic closed-loop grasping phase of tasks pouring and reorient+unscrew, where each object is manually disturbed from one to three times. The red arrow indicates the direction of the manually moved object (interfering). The cyan arrow and yellow arrow indicate the movement direction of the left and right robotic arms (chasing) respectively. Best to view after zooming in.
Table 2: Quantitative comparison results of success rates for the few-shot BiDemoSyn under in-distribution and out-of-distribution evaluations. The # means the BiDemoSyn training data size. Although the number of demonstrations used for training is always the same for each setting, the number of seed demonstrations used for trajectory synthesis is different.
  in-distribution evaluations out-of-distribution evaluations
Trained
Policy
Few-Shot
BiDemoSyn
plugpen
inserting
unscrew
pouring
pressing
reorient
Average
Success
Rate
plugpen
inserting
unscrew
pouring
pressing
reorient
Average
Success
Rate
#3888 #7776 #1152 #2592 #1296 #1008 #2916 #3888 #1008 #972 #324 #756
  EquiBot 1-shot 28/30 28/30 26/30 25/30 24/30 25/30 86.7% 18/30 20/30 24/30 21/30 17/30 20/30 66.7%
5-shot 28/30 28/30 26/30 26/30 25/30 25/30 87.8% 20/30 21/30 24/30 23/30 18/30 21/30 70.0%
10-shot 28/30 28/30 26/30 27/30 25/30 26/30 88.9% 21/30 21/30 25/30 23/30 18/30 22/30 71.7%
15-shot 28/30 28/30 27/30 27/30 26/30 26/30 90.0% 21/30 22/30 25/30 23/30 18/30 22/30 72.2%
20-shot 28/30 28/30 27/30 27/30 26/30 26/30 90.0% 21/30 22/30 25/30 23/30 18/30 22/30 72.2%
D.4 Few-Shot-Based Demonstration Synthesis
While BiDemoSyn is designed around a one-shot paradigm, the framework naturally extends to a few-shot setting, where multiple real demonstrations may serve as complementary seeds. This subsection studies the effect of incorporating a small number of additional demonstrations (5–20 samples) obtained through standard teleoperation or kinesthetic teaching (used in this study).
Few-Shot Integration into BiDemoSyn. Given multiple seed demonstrations, we incorporate them in two ways: (1) Enhanced Object Variation Coverage. Additional demonstrations expand the diversity of observed object geometries, grasp points, and intermediate contact poses. These are incorporated into the variant blocks during trajectory adaptation. (2) Improved Robustness of Reference Motion. Multiple demonstrations allow BiDemoSyn to aggregate or select more reliable canonical motion patterns for the invariant blocks, mitigating single-demonstration idiosyncrasies. Under these two settings, the followed overall synthesis pipeline remains unchanged.
Experimental Findings. We have summarized results of this ablation study in Tab. 2. These new experiments show: (1) In-distribution (ID) performance. Adding few-shot data provides little to no improvement. Policies trained with BiDemoSyn’s one-shot–generated datasets already exhibit near-saturation performance for seen objects and placements. (2) Out-of-distribution (OOD) performance. Few-shot data provides moderate gains, primarily by increasing the number of observed object instances and geometric variations, thereby improving generalization to unseen shapes and dimensions. These findings align with the intuition that BiDemoSyn already produces high-quality, physically faithful demonstrations from a single demonstration (refer Fig. 5). Additional few-shot demonstrations mainly help by expanding the representational coverage rather than correcting failure cases caused by hardware limitations of the embodied robot entity itself.
Additional Conclusion. The few-shot extension confirms that BiDemoSyn supports scalable integration of multiple demonstrations, maintains strong performance in the one-shot setting, benefits from few-shot seeds mostly in the OOD regime, and requires no modification to its core synthesis pipeline. This suggests a promising avenue for future research: exploring hybrid one-shot + few-shot paradigms that balance data-collection cost with maximal generalization capability.
D.5 VFM-based Object Anchoring in Clutter
To evaluate the applicability of BiDemoSyn under more realistic and visually complex environments, we additionally examine its object-anchoring capability when the target objects are placed in cluttered scenes with distractors, varied background patterns, and partial occlusions. Although the main experiments intentionally isolate environmental variables to study the core challenge of real-world demonstration synthesis, BiDemoSyn is fully compatible with more complex perception conditions due to its reliance on Vision Foundation Models (VFMs) for object segmentation.
Clutter-Aware Object Localization. We employ Florence-2 xiao2024florence and SAM2 ravi2024sam to detect and segment the task-relevant items (e.g., bottle, mug) using either text queries or exemplar-based prompts. In cluttered tabletop scenes, the VFM stack consistently produces high-quality masks, even when distractors share similar colors or shapes. Qualitatively, the predicted masks remain stable across moderate occlusions, enabling the system to isolate the target object with minimal ambiguity. Fig. 17 illustrates representative examples where both bottle and mug are correctly segmented despite surrounding tools, utensils, packaging, and nonrigid paper balls.
Implications for Demonstration Synthesis. Because BiDemoSyn’s variant-block adaptation depends only on (1) reliable 2D masks and (2) centroid- and principal-axis–based orientation inference, accurate object anchoring is the only requirement for extending the system to cluttered environments. Our experiments show that the VFM-based perception pipeline provides sufficiently robust segmentation to support (a) pose-aligned trajectory modulation and (b) collision-aware grasp adjustments, even when the background is visually nonuniform.
Limitations and Future Directions. While VFMs significantly enhance robustness to visual complexity, extreme occlusions or heavy inter-object entanglement may still degrade performance. Extending BiDemoSyn to handle these cases will require multi-view sensing, active view planning, or integrating VLM-driven object-query disambiguation. Nevertheless, these initial results demonstrate that VFM-based anchoring provides a practical and reliable foundation for deploying BiDemoSyn beyond clean tabletop settings, thereby enabling future extensions to household, warehouse, and mobile manipulation scenarios.
Figure 17: Top row: Example cluttered tabletop scenes used to evaluate the perception front-end of BiDemoSyn. Each scene includes the target objects for the pouring task—two bottles (text prompts are blue bottle and green bottle) and two mugs (text prompts blue cup and green cup)—surrounded by visually distracting household items. Two additional tool-like distractors (brush, dustpan) are intentionally introduced to support future extensions such as dual-arm sweeping tasks. Bottom row: High-quality 2D object masks produced by Florence-2 + SAM2 for all queried categories. Despite clutter, color similarity, and semantic ambiguity, the VFM stack consistently isolates each target object with precise boundaries, providing reliable anchors for subsequent pose inference and variant-block adaptation in BiDemoSyn.
Appendix E Summary, Reflection and Outlook
This work presents BiDemoSyn, a framework that bridges the reality gap in bimanual imitation learning by synthesizing diverse, physically feasible demonstrations from a single human example. Through systematic decomposition, vision-guided alignment, and hierarchical optimization, our method eliminates simulation dependencies while scaling data collection by orders of magnitude. Experiments across six contact-rich tasks validate its efficiency (5 seconds per demo), scalability, and generalization to unseen spatial and geometric variations.
However, limitations persist: BiDemoSyn currently assumes quasi-static motions, struggles with highly dynamic tasks (e.g., catching), and faces challenges in extreme shape variations (e.g., deformable objects). Failures primarily stem from orientation estimation errors and grasp-time collisions, highlighting needs for category-agnostic pose estimators and contact-aware motion planners. Future work will extend BiDemoSyn to deformable object manipulation, integrate dynamic perception for moving targets, and explore hybrid learning paradigms (e.g., combining synthetic demos with reinforcement learning). By democratizing real-world demonstration synthesis, this work aims to accelerate progress toward generalizable robotic manipulation.
Appendix F Use of Large Language Models
In preparing this manuscript, we employed the GPT-4o language model strictly for text editing purposes, such as improving readability and polishing grammar. The model was not used to generate research ideas, design methodologies, conduct experiments, or analyze results. All substantive contributions remain entirely those of the authors.

### v2

@@
 This enables the generation of thousands of diverse and feasible demonstrations within several hour, without repeated teleoperation or reliance on imperfect simulation.
-Across six dual-arm tasks, we show that policies trained on BiDemoSyn data generalize robustly to novel object poses and shapes, significantly outperforming recent baselines.
+Across six dual-arm tasks, we show that policies trained on BiDemoSyn data generalize robustly to novel object poses and shapes, significantly outperforming recent strong baselines.
+Beyond the one-shot setting, BiDemoSyn naturally extends to few-shot-based synthesis, improving object-level diversity and out-of-distribution generalization while maintaining strong data efficiency.
+Moreover, policies trained on BiDemoSyn data exhibit zero-shot cross-embodiment transfer to new robotic platforms, enabled by object-centric observations and a simplified 6-DoF end-effector action representation that decouples policies from embodiment-specific dynamics.
 By bridging the gap between efficiency and real-world fidelity, BiDemoSyn provides a scalable path toward practical imitation learning for complex bimanual manipulation without compromising physical grounding.
-1 Introduction
+I Introduction
 Recent advances in robot manipulation have been propelled by data-driven imitation learning, where visuomotor policies trained on large-scale demonstration datasets enable dexterous single-arm and bimanual tasks previously deemed intractable.
-Methods like Action Chunking Transformer (ACT) zhao2023learning, Diffusion Policy (DP) chi2023diffusion and $\pi_{0}$ black2024pi0 exemplify this trend, achieving remarkable performance in contact-rich scenarios such as continuously pushing, battery assembly, and coordinated garment folding.
+Methods like ACT [103], DP [16] and $\pi_{0}$ [5] exemplify this trend, achieving remarkable performance in contact-rich scenarios such as continuously pushing, battery assembly, and coordinated garment folding.
 These successes, however, hinge on a critical yet underappreciated premise: the availability of diverse, high-quality demonstrations, which are typically collected via expert teleoperation.
 As the community shifts toward more complex long-horizon bimanual tasks, the data scalability bottleneck becomes increasingly apparent.
-Even those largest manipulation datasets like RH20T fang2024rh20t, DROID khazatsky2024droid, Open X-Embodiment o2024open and AgiBot World bu2025agibot struggle to provide sufficient geometric and kinematic diversity for robust generalization, underscoring a pivotal question: how to scale real-world demonstration collection without compromising physical fidelity?
-[…]
+Even those largest manipulation datasets like RH20T [22], DROID [44], Open X-Embodiment [68] and AgiBot World [6] struggle to provide sufficient geometric and kinematic diversity for robust generalization, underscoring a pivotal question: how to scale real-world bimanual demonstration collection without compromising physical fidelity?
 The answer, unfortunately, lies in a fundamental trade-off.
 Existing data acquisition methods fall into two paradigms, each with critical shortcomings.
-On one hand, human-operated systems (e.g., the ALOHA series zhao2023learning; fu2024mobile; zhao2024aloha) yield physically-accurate trajectories but demand prohibitive expertise, rendering them impractical for scaling to long-horizon tasks.
-On the other hand, simulation-based methods (e.g., MimicGen mandlekar2023mimicgen, RoboGen wang2024robogen, RoboCasa nasiriany2024robocasa, and DexMimicGen jiang2024dexmimicgen) bypass human labor through programmatic data generation utilizing massive 3D assets, yet discrepancies in contact dynamics and visual rendering inevitably plague real-world deployment bharadhwaj2024position; mccarthy2024towards.
+On one hand, human-operated systems (e.g., the ALOHA series [103, 26, 104]) yield physically-accurate trajectories but demand prohibitive expertise, rendering them impractical for scaling to long-horizon tasks.
+On the other hand, simulation-based methods (e.g., MimicGen [62, 40], RoboGen [85], and RoboCasa [66]) bypass human labor through programmatic data generation utilizing massive 3D assets, yet discrepancies in contact dynamics and visual rendering inevitably plague real-world deployment [3, 64].
 Thus, the tension between data quality (grounded in reality) and quantity (enabled by automation) persists as a key barrier to scalable imitation learning, particularly for contact-rich bimanual tasks requiring precise dual-arm coordination.
-To bridge this gap, we present BiDemoSyn that synthesizes real-world bimanual demonstrations from a single exemplar, by algorithmically amplifying task semantics through a hierarchical decomposition of invariance and adaptability.
-Our key insight lies in recognizing that complex manipulation tasks exhibit a duality: while certain motion primitives (e.g., stable grasp sequences, dual-arm synchronization) remain invariant across instances, others (e.g., trajectory segments adapting to object shifting) require dynamic adjustments to contextual variations.
-Through a three-stage process, BiDemoSyn first deconstructs the initial demonstration into invariant primitives and geometry-sensitive components, then leverages visual perception to spatially align these elements with novel scenes, enabling generalization across object geometries and workspace layouts with minor human effort (primarily involving reinitializing manipulated objects).
+To bridge this gap, we present BiDemoSyn that synthesizes real-world bimanual demonstrations from a single exemplar, by algorithmically amplifying task semantics through a hierarchical decomposition of invariance and adaptability (refer Fig.
+LABEL:BiDemoSyn and Fig.
+1).
+Our motivation is inherited from the well-established single-arm one-shot imitation learning [63, 60, 4, 100, 21, 106].
+The key insight lies in recognizing that complex manipulation tasks exhibit a duality: while certain motion primitives (e.g., stable grasp sequences, dual-arm synchronization) remain invariant across instances, others (e.g., trajectory segments adapting to object shifts) require dynamic adjustments to contextual variations.
+Specifically, through a three-stage process, BiDemoSyn first deconstructs the initial demonstration into invariant primitives and geometry-sensitive components.
+Then, it leverages visual perception to spatially align these elements in novel scenes, enabling generalization across object geometries and workspace layouts with minor effort (primarily involving reinitializing objects).
 Finally, a physics-aware optimizer jointly modulates dual-arm trajectories, injecting diversity into adaptable components while enforcing real-world constraints on collision avoidance and kinematic coordination.
 This integrated approach achieves One-to-Many demonstration synthesis: a single exemplar spawns hundreds of trajectories that preserve task intent while adapting to unseen configurations, all grounded in real-world dynamics rather than simulated proxies.
-Examples are shown in Fig.
-1 and Fig.
-2.
-Unlike prior simulation-based methods mandlekar2023mimicgen; wang2024robogen; nasiriany2024robocasa; jiang2024dexmimicgen, BiDemoSyn operates entirely within the physical domain, ensuring synthesized data inherits the fidelity of human demonstrations.
+Unlike prior simulation-based methods [62, 85, 66, 40], BiDemoSyn operates entirely in the physical domain, ensuring synthesized data inherits the fidelity of human demonstrations.
 By unifying algorithmic scalability with physical realism, our framework empowers policies to transcend the limitations of narrow manually-collected data distributions.
 Experiments on real dual-arm platforms validate the effectiveness of BiDemoSyn.
-And visuomotor policies trained with our synthesized demonstrations can achieve superior success rates and generalization in tasks requiring precise contact coordination (e.g., small hole insertion, rotation angle control, purposeful handover and articulated object manipulation).
+Visuomotor policies trained with synthesized demonstrations can achieve superior success rates and generalization in tasks requiring precise contact coordination (e.g., small hole insertion, rotation angle control, purposeful handover and articulated object manipulation).
 Our contributions are threefold: (1) One-Shot Synthesis Framework: A systematic pipeline combining task decomposition, vision-guided adaptation, and contact-aware trajectory optimization to generate scalable real-world bimanual demonstrations.
@@
 (3) Empirical Validation in Complex Tasks: Comprehensive real-robot experiments demonstrating significant improvements in policy robustness and cross-configuration generalization on various bimanual manipulation tasks.
-2 Related Works
+II Related Works
 Bimanual Robotic Manipulation.
-Prior literatures primarily target task-specific challenges (such as cloth folding colome2018dimensionality; weng2022fabricflownet; canberk2023cloth, untangling grannen2021untangling; peng2024tiebot, untwisting lin2024twisting; bahety2024screwmimic, and handover huang2023dynamic; li2023efficient) through handcrafted controllers or narrow skill libraries, limiting generalization due to rigid motion primitives and excessive assumptions.
-General methods xie2020deep; chitnis2020efficient; chen2022towards; chen2023bi; hartmann2022long; zhaodual2023afford either rely on predefined hierarchical arm roles (e.g., leader-follower krebs2022bimanual; liu2022robot; grotz2024peract2 and stabilizer-actor grannen2023stabilize; liu2024voxact) or hinge generalization entirely on dataset diversity, both of which are ultimately bottlenecked by the impractical costs of data acquisition.
-Recently, ALOHA series zhao2023learning; fu2024mobile; zhao2024aloha have revolutionized bimanual manipulation by dexterous low-cost teleoperation.
-Meanwhile, the visuomotor policy learning paradigm chi2023diffusion; ze2024dp3; yang2024equibot based on diffusion models ho2020denoising; song2021denoising has largely expanded the action modeling space and data utilization efficiency.
-Their followers team2024octo; kim2024openvla; liu2025rdt; black2024pi0; pertsch2025fast; lin2025data expect to train universal policies using large-scale teleoperation data but face scalability barriers from human effort.
-To improve reachability and dexterity, some studies use specialized hardware (e.g., multi-finger hands wang2024dexcap; shaw2024bimanual; fu2024humanplus; cheng2024open or tactile sensors lin2024learning; chen2024arcap), which complicate real-world adoption.
+Prior literatures primarily target task-specific challenges (such as cloth folding [20, 88, 7], untangling [29, 70], untwisting [54, 1], and handover [37, 51]) through handcrafted controllers or narrow skill libraries, limiting generalization due to rigid motion primitives and excessive assumptions.
+General methods [90, 18, 12, 10, 33, 105] either rely on predefined hierarchical arm roles (e.g., leader-follower [48, 57, 32] and stabilizer-actor [30, 56]) or hinge generalization entirely on dataset diversity, both of which are ultimately bottlenecked by the impractical costs of data acquisition.
+Recently, ALOHA series [103, 26, 104] have revolutionized bimanual manipulation by dexterous low-cost teleoperation.
+Meanwhile, the visuomotor policy learning paradigm [16, 98, 94] based on diffusion models [35, 79] has largely expanded the action modeling space and data utilization efficiency.
+Their followers [80, 45, 58, 5, 71, 53] expect to train universal policies using large-scale teleoperation data but face scalability barriers from human effort.
+To improve reachability and dexterity, some studies use specialized hardware (e.g., multi-finger hands [82, 76, 25, 15] or tactile sensors [55, 9]), which complicate real-world adoption.
 In contrast, we utilize a fixed-base dual-arm platforrm with parallel grippers, synthesizing diverse demonstrations from a single example.
@@
 Bimanual Demonstration Collection.
-The dominant path for acquiring bimanual demonstrations is human teleoperation khazatsky2024droid; bu2025agibot, which delivers high-quality data but suffers from prohibitive scalability costs.
-To alleviate it, two alternative routes have emerged: simulation-based synthesis garrett2024skillmimicgen; hua2024gensim2; liang2024make; yang2024physcene; wang2024gensim and learning from human videos ponimatkin20256d; ye2025video2policy; zhao2025taste; kareer2024egomimic; bharadhwaj2024track2act.
-The former, exemplified by MimicGen mandlekar2023mimicgen and DexMimicGen jiang2024dexmimicgen, augments a few demonstrations by generating variations in simulation using geometric transformations and kinematic constraints.
-Analogously, RoboGen wang2024robogen and RoboCasa nasiriany2024robocasa leverage 3D assets to procedurally generate full-scene manipulation data, yet such methods inevitably inherit sim-to-real gaps, from unrealistic textures to unphysical dynamics.
-The latter extracts bimanual hand-object manipulation from egocentric videos zhan2024oakink2; liu2024taco; grauman2024ego and maps them to dual arms via motion retargeting li2024okami; kerr2024robot; chen2024object; chen2024vividex or non-privileged representations (e.g., keypoints papagiannis2024rx; gao2024bi; wen2023any, affordances ju2024robo; nasiriany2024rt and correspondences peng2024tiebot; ko2024learning; zhang2024one).
+The dominant path for acquiring bimanual demonstrations is human teleoperation [44, 6], which delivers high-quality data but suffers from prohibitive scalability costs.
+To alleviate it, two alternative routes have emerged: simulation-based synthesis [28, 36, 52, 96, 83] and learning from human videos [72, 97, 102, 42, 2].
+The former, exemplified by MimicGen [62] and DexMimicGen [40], augments a few demonstrations by generating variations in simulation using geometric transformations and kinematic constraints.
+Analogously, RoboGen [85] and RoboCasa [66] leverage 3D assets to procedurally generate full-scene manipulation data, yet such methods inevitably inherit sim-to-real gaps, from unrealistic textures to unphysical dynamics.
+The latter extracts bimanual hand-object manipulation from egocentric videos [99, 59, 31] and maps them to dual arms via motion retargeting [50, 43, 11, 13] or non-privileged representations (e.g., keypoints [69, 27, 87], affordances [41, 65] and correspondences [70, 46, 101]).
 While human videos are abundant and low-cost, large morphological mismatches between humanity and robotic embodiments often need heuristic translation rules and hinder direct applicability.
@@
 Several concurrent works exhibit notable limitations.
-DemoGen xue2025demogen and YOTO zhou2025you diversify demonstrations via 3D editing of the initial single-view point cloud, but perspective ambiguities induce visual artifacts in synthesized data.
-ODIL wang2025one avoids editing by segmenting objects for visual servoing and iteratively replanning trajectories, which is cumbersome for scalability.
-MoMaGen li2025momagen is a long-horizon mobile+bimanual demonstrations generator yet mainly focusing on the simulation domain.
+DemoGen [93] and YOTO [107] diversify demonstrations via 3D editing of the initial single-view point cloud, but perspective ambiguities induce visual artifacts in synthesized data.
+ODIL [84] avoids editing by segmenting objects for visual servoing and iteratively replanning trajectories, which is cumbersome for scalability.
+MoMaGen [49] is a long-horizon mobile bimanual demonstrations generator yet mainly focusing on the simulation.
 Our BiDemoSyn sidesteps these issues: synthesizing trajectories directly in real-world ensures visual-physical consistency without manual edits or robot replaying.
-3 Preliminaries
+III Preliminaries
 Problem Formulation.
 Bimanual imitation learning aims to train visuomotor policies $\pi_{\theta}(\textbf{o}_{t})\rightarrow\textbf{a}_{t}$, where $\textbf{o}_{t}$ denotes multimodal observations (e.g., RGB images, joint states or point clouds) and $\textbf{a}_{t}$ represents dual-arm actions.
-While architectures like ACT zhao2023learning and DP chi2023diffusion enhance policy expressivity, their generalization critically depends on expert demonstrations $\mathcal{D}\!=\!\{\tau_{i}\}_{i=0}^{N}$ that densely span the feasible state-action manifold.
+While architectures like ACT [103] and DP [16] enhance policy expressivity, their generalization critically depends on expert demonstrations $\mathcal{D}\!=\!\{\tau_{i}\}_{i=0}^{N}$ that densely span the feasible state-action manifold.
 Existing approaches either collect $\mathcal{D}$ via labor-intensive teleoperation or generate data in simulation, both failing to balance real-world fidelity and scalability.
@@
 Our $\pi_{\theta}$ maps observations $\textbf{o}_{t}$ to dual-arm actions $\textbf{a}_{t}\!=\!\left[\textbf{a}_{t}^{L},\textbf{a}_{t}^{R}\right]\!\in\!\mathbb{R}^{2d}$, where $\textbf{a}_{t}^{L}$, $\textbf{a}_{t}^{R}$ denotes left/right end-effector poses.
-We adapt a diffusion policy model chi2023diffusion; ze2024dp3; yang2024equibot to generate action sequences $\textbf{A}\!=\!\left[\textbf{a}_{t:{t+H}}^{L},\textbf{a}_{t:{t+H}}^{R}\right]\!\in\!\mathbb{R}^{2H\times d}$ over a horizon $H$.
+We adapt a diffusion policy model [16, 98, 94] to generate action sequences $\textbf{A}\!=\!\left[\textbf{a}_{t:{t+H}}^{L},\textbf{a}_{t:{t+H}}^{R}\right]\!\in\!\mathbb{R}^{2H\times d}$ over a horizon $H$.
 The diffusion process iteratively denoises a noisy action sequence $\textbf{A}_{k}$ toward kinematic feasibility over $K$ steps:
 $\centering\textbf{A}_{k}=\sqrt{\alpha_{k}}\textbf{A}_{0}+\sqrt{1-\alpha_{k}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{\Sigma}),\@add@centering$ (1)
-where $\alpha_{k}$ follows a cosine schedule nichol2021improved.
+where $\alpha_{k}$ follows a cosine schedule [67].
 And $\mathbf{\Sigma}\!\in\!\mathbb{R}^{2H\times 2H}$ is a block-diagonal covariance matrix $\left[\begin{array}[]{cc}\mathbf{\Sigma}_{L}&\rho\mathbf{\Sigma}_{LR}\\ \rho\mathbf{\Sigma}^{\top}_{LR}&\mathbf{\Sigma}_{R}\end{array}\right]$ encoding arm coordination priors, where $\mathbf{\Sigma}_{L}$, $\mathbf{\Sigma}_{R}$ governing temporal smoothness per arm, $\mathbf{\Sigma}_{LR}$ modeling inter-arm dependencies, and $\rho\!\in\![0,1]$ controlling coordination strength.
@@
 $\centering\mu_{\theta}(\textbf{A}_{k},\textbf{o}_{t},k)=\texttt{UNet}_{\theta}(\texttt{Concat}\left[\textbf{A}_{k},\mathsf{E}(\textbf{o}_{t})\right],k),\@add@centering$ (2)
-where $\mathsf{E}(\textbf{o}_{t})$ is a vision encoder (ResNet he2016deep or PointNet qi2017pointnet) extracting latent features.
+where $\mathsf{E}(\textbf{o}_{t})$ is a vision encoder (ResNet [34] or PointNet [73]) extracting latent features.
 The $\texttt{UNet}_{\theta}$ employs asymmetric layers with early stages processing each arm independently, while deeper layers fusing bilateral coordinated features.
@@
 Then, we can apply our method to complete data collection efficiently and conveniently in real-world.
-It is best to zoom in to view the details.
-4 Methodology
-Here, we detail BiDemoSyn for synthesizing bimanual demonstrations $\mathcal{D}_{syn}$ from a single exemplar $\tau$ (teleoperated or kinesthetically guided) with three stages: Sec.
-4.1 Deconstruction of One-Shot Teaching, which extracts invariant patterns and adaptable primitives; Sec.
-4.2 Vision-Based Initial Frame Alignment, enabling generalization across geometric variations via efficient scene perception; and Sec.
-4.3 Trajectory Modulation and Optimization, ensuring physical feasibility through hierarchical kinematic constraints.
+IV Methodology
+Here, we detail BiDemoSyn for synthesizing bimanual demonstrations $\mathcal{D}_{syn}$ from a single exemplar $\tau$ with three compact stages (see Fig.
+1): Sec.
+IV-A Deconstruction of One-Shot Teaching, which extracts invariant patterns and adaptable primitives; Sec.
+IV-B Vision-based Initial Frame Alignment, enabling generalization across geometric variations via efficient scene perception; and Sec.
+IV-C Trajectory Modulation and Optimization, ensuring physical feasibility via hierarchical kinematic constraints.
 Each stage sequentially addresses scalability, adaptability, and real-world fidelity to bridge the gap between data efficiency and policy robustness.
-The overall framework is shown and illustrated in Fig.
-2.
-4.1 Deconstruction of One-Shot Teaching
-Given a single demonstration $\tau\!=\!\{\textbf{o}_{t},\textbf{a}_{t}\}_{t=0}^{T}$, we decompose it into a bimanual execution blocks set $\{\mathcal{B}_{i}\}_{i=1}^{n}$, where each block $\mathcal{B}_{i}\!=\!(\textbf{s}_{i}^{L},\textbf{s}_{i}^{R})$ represents a discrete phase of dual-arm interaction.
-Here, $\textbf{s}_{i}^{L},\textbf{s}_{i}^{R}\!\in\!\mathbb{R}^{d}$ denote the left/right arm state sequences (including end-effector 6-DoF poses and gripper status) within block $\mathcal{B}_{i}$.
+IV-A Deconstruction of One-Shot Teaching
+Given a demonstration $\tau\!=\!\{\textbf{o}_{t},\textbf{a}_{t}\}_{t=0}^{T}$, we decompose it into a bimanual execution blocks set $\{\mathcal{B}_{i}\}_{i=1}^{n}$, where each block $\mathcal{B}_{i}\!=\!(\textbf{s}_{i}^{L},\textbf{s}_{i}^{R})$ represents a discrete phase of dual-arm interaction.
+Here, $\textbf{s}_{i}^{L},\textbf{s}_{i}^{R}\!\in\!\mathbb{R}^{d}$ denote the left/right arm state sequences (e.g., end-effector 6-DoF poses and gripper status) within block $\mathcal{B}_{i}$.
 Blocks are categorized based on motion coordination.
-For the single-arm motion, one arm exhibits significant state transitions while the other remains static.
+For single-arm motion, one arm exhibits significant transitions while the other remains static.
 Formally, for threshold $\delta$ (minimal motion saliency) and $\zeta$ (static tolerance):
-$\centering(\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\leq\!\zeta)\;\vee\;(\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\leq\!\zeta).\@add@centering$ (4)
-While, for the dual-arm coordination, both arms synchronously or asynchronously adjust states:
+$(\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\leq\!\zeta)\;\vee\;$ (4)
+$(\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\leq\!\zeta).$
+While, for dual-arm coordination, both arms synchronously or asynchronously adjust states:
 $\centering\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\geq\!\delta.\@add@centering$ (5)
-From Blocks to Atomic Execution Primitives (AEPs).
+IV-A1 From Blocks to Atomic Execution Primitives (AEPs)
 To enable modular adaptation, each block $\mathcal{B}_{i}$ is further refined into a minimal motion unit AEP, where an arm undergoes a salient state transition (e.g., gripper closure, moving or pose shifts).
 For arm $A\!\in\!\{L,R\}$, an AEP is defined as:
-$\centering\texttt{AEP}_{A}=\left\{\textbf{s}_{t}^{A}\rightarrow\textbf{s}_{t+\Delta t}^{A}\;|\;\left\|\textbf{s}_{t+\Delta t}^{A}-\textbf{s}_{t}^{A}\right\|_{2}\geq\gamma,\Delta t\leq T_{\textrm{max}},\texttt{CollisionFree}(\textbf{s}_{[t,\Delta t]}^{A})\right\},\@add@centering$ (6)
-where $\gamma$ is a distance threshold, $T_{\textrm{max}}$ limits execution duration, and $\texttt{CollisionFree}(\cdot)$ enforces no collisions between the arm and objects/scene (optimized via subsequent forward kinematics).
-Unlike keyframe-based segmentation james2022coarse; shridhar2023perceiver; ma2024hierarchical, AEPs ignore acceleration profiles (assuming quasi-static motions) and prioritize task-oriented transitions over temporal granularity.
-Semantic Categorization for Adaptive Synthesis.
+$\centering\texttt{AEP}_{A}=$ $\{\textbf{s}_{t}^{A}\rightarrow\textbf{s}_{t+\Delta t}^{A}\;|\;\left\|\textbf{s}_{t+\Delta t}^{A}-\textbf{s}_{t}^{A}\right\|_{2}\geq\gamma,$ (6)
+$\Delta t\leq T_{\textrm{\scriptsize max}},\texttt{CollisionFree}(\textbf{s}_{[t,\Delta t]}^{A})\},$
+where $\gamma$ is a distance threshold, $T_{\textrm{\scriptsize max}}$ limits execution duration, and $\texttt{CollisionFree}(\cdot)$ enforces no collisions between the arm and objects/scene (optimized via forward kinematics).
+Unlike keyframe-based segments [39, 78, 61, 107], AEPs ignore acceleration profiles (assuming quasi-static motions) and prioritize task-oriented transitions over temporal granularity.
+IV-A2 Semantic Categorization for Adaptive Synthesis
 The final step categorizes refined blocks into invariant and variable types.
@@
 The variable blocks, such as object-centric goal-conditioned grasping, adapt to instance-level geometric variations including object poses and shapes through:
-$\centering\mathcal{B}^{\prime}_{i}=\mathcal{B}_{i}\circ\Phi(g,\textbf{o}_{\textrm{novel}}),\@add@centering$ (7)
-where $\Phi(\cdot)$ is a vision-driven adapter, $g$ is the task goal, and $\textbf{o}_{\textrm{novel}}$ is the novel observation.
-This structured decomposition enables selective diversification, allowing variable blocks to adapt to new scenarios while preserving task fidelity in invariant blocks.
-4.2 Vision-based Initial Frame Alignment
+$\centering\mathcal{B}^{\prime}_{i}=\mathcal{B}_{i}\circ\Phi(g,\textbf{o}_{\textrm{\scriptsize novel}}),\@add@centering$ (7)
+where $\Phi(\cdot)$ is a visual adapter, $g$ is the task goal, and $\textbf{o}_{\textrm{\scriptsize novel}}$ is the novel observation.
+This structured decomposition enables selective diversification, allowing variable blocks to adapt to new scenes while preserving task fidelity in invariant blocks.
+IV-B Vision-based Initial Frame Alignment
 The variable blocks identified in Sec.
-4.1 (e.g., initial grasping) require adapting to novel object poses and geometries.
+IV-A require adapting to novel object poses and geometries.
 To achieve this, we propose a vision-driven adapter $\Phi(\cdot)$ that generalizes object-centric interactions in the given exemplar to new scenes.
-Given a novel observation $\textbf{o}_{\textrm{novel}}$, our method executes three steps: object perception, state estimation, and pose alignment, ensuring the initial robotic grasp aligns with the task goal $g$.
+Given a novel observation $\textbf{o}_{\textrm{\scriptsize novel}}$, our method executes three steps: object perception, state estimation, and pose alignment, ensuring the initial robotic grasp aligns with the task goal $g$.
 Illustrations are shown in Fig.
-3.
-Figure 3: Illustrations of the initial frame alignment stage applied to tasks pouring (left and middle) and reorient (right).
+2.
+Figure 2: Illustrations of the initial frame alignment applied to tasks pouring (left and middle) and reorient (right).
 It shows that we can automatically adjust the grasp pose after the position, orientation and shape of the manipulated object changes.
-Object Perception.
-We first detect and segment the target object in $\textbf{o}_{\textrm{novel}}$ using an open-vocabulary detector (e.g., YOLO-World cheng2024yolo or YOLOE wang2025yoloe) or vision foundation models (e.g., Florence2 xiao2024florence with SAM2 ravi2024sam for robustness to rare categories).
+IV-B1 Object Perception
+We first detect and segment the target object in $\textbf{o}_{\textrm{\scriptsize novel}}$ using an open-vocabulary detector (e.g., YOLO-World [14] or YOLOE [81]) or vision foundation models (e.g., Florence2 [89] with SAM2 [74] for robustness to rare categories).
 In practice, we prioritize foundation models to avoid detection failures.
-Now let $M\!\subseteq\!\textbf{o}_{\textrm{novel}}$ denote the obtained object’s binary mask.
-State Estimation.
-Instead of uitlizing category-level CAD-based 6D pose estimation models (e.g., FoundationPose wen2024foundationpose), we estimate the instance-level 6D object pose $\mathbf{P}_{\textrm{obj}}\!\in\!SE(3)$ using geometry-aware processing.
-Sepcifically, we adopt classic image moments chaumette2004image; kotoulas2007accurate to calculate the object mask centroid c and extract principal axes R:
+Now let $M\!\subseteq\!\textbf{o}_{\textrm{\scriptsize novel}}$ denote the obtained object’s binary mask.
+IV-B2 State Estimation
+Instead of uitlizing category-level CAD-based 6D pose estimation models (e.g., FoundationPose [86]), we estimate the instance-level 6D object pose $\mathbf{P}_{\textrm{\scriptsize obj}}\!\in\!SE(3)$ using geometry-aware processing.
+Sepcifically, we adopt classic image moments [8, 47] to calculate the object mask centroid c and extract principal axes R:
 $\centering\left\{\begin{array}[]{ccl}\textbf{c}&=&\frac{1}{|M|}\sum_{(u,v)\in M}(u,v,\textrm{d}(u,v))\\ \textbf{R}&=&\texttt{PCA}(\{(u,v,\textrm{d}(u,v))\;|\;(u,v)\in M\})\end{array}\right$.\@add@centering (8)
@@
 PCA is used to fit 3D points of $M$ to determine orientation $\textbf{R}\!\in\!SO(3)$.
-This yields $\mathbf{P}_{\textrm{obj}}\!=\!(\textbf{R},\textbf{c})$, robust to arbitrary object states (e.g., fallen, inverted).
-Pose Alignment.
-Let $\mathbf{P}_{\textrm{demo}}$ denote the object pose in the given demonstration.
-We compute the rigid transformation $\mathbf{T}\!\in\!SE(3)$ that maps $\mathbf{P}_{\textrm{demo}}$ to $\mathbf{P}_{\textrm{obj}}$.
-This transformation is applied to the initial grasp pose $\mathbf{P}_{\textrm{grasp,demo}}$ in the variable block $\mathcal{B}_{i}$, yielding the adapted grasp pose:
-$\centering\mathbf{P}_{\textrm{grasp,novel}}=\mathbf{T}\cdot\mathbf{P}_{\textrm{grasp,demo}}=(\mathbf{P}_{\textrm{obj}}\cdot\mathbf{P}^{-1}_{\textrm{demo}})\cdot\mathbf{P}_{\textrm{grasp,demo}}.\@add@centering$ (9)
-The robot arm then executes a collision-free motion to reach $\mathbf{P}_{\textrm{grasp,novel}}$, ensuring goal-conditioned grasping without relying on task-irrelevant dense grasp proposals produced via 6-DoF grasp pose detectors fang2020graspnet; fang2023anygrasp.
-By integrating with the decomposition stage, the adapted grasp pose directly modifies the variable block $\mathcal{B}_{i}$ into $\mathcal{B}^{\prime}_{i}$, enabling synthesis of new trajectories.
-4.3 Trajectory Modulation and Optimization
-After deconstructing the demonstration into blocks (Sec.
-4.1) and adapting variable blocks via vision-based alignment (Sec.
-4.2), the synthesized trajectory requires two critical refinements to ensure executability and robustness as below:
-Collision-Aware Validation.
+This yields $\mathbf{P}_{\textrm{\scriptsize obj}}\!=\!(\textbf{R},\textbf{c})$, robust to arbitrary object states (e.g., fallen, inverted).
+IV-B3 Pose Alignment
+Let $\mathbf{P}_{\textrm{\scriptsize demo}}$ denote the object pose in the given demonstration.
+We compute the rigid transformation $\mathbf{T}\!\in\!SE(3)$ that maps $\mathbf{P}_{\textrm{\scriptsize demo}}$ to $\mathbf{P}_{\textrm{\scriptsize obj}}$.
+This transformation is applied to the initial grasp pose $\mathbf{P}_{\textrm{\scriptsize grasp,demo}}$ in the variable block $\mathcal{B}_{i}$, yielding the adapted grasp pose:
+$\centering\mathbf{P}_{\textrm{\scriptsize grasp,novel}}=\mathbf{T}\cdot\mathbf{P}_{\textrm{\scriptsize grasp,demo}}=(\mathbf{P}_{\textrm{\scriptsize obj}}\cdot\mathbf{P}^{-1}_{\textrm{\scriptsize demo}})\cdot\mathbf{P}_{\textrm{\scriptsize grasp,demo}}.\@add@centering$ (9)
+The robot arm then executes a collision-free motion trajectory to reach $\mathbf{P}_{\textrm{\scriptsize grasp,novel}}$, ensuring goal-conditioned grasping without relying on task-irrelevant dense grasp proposals produced via 6-DoF grasp pose detectors [24, 23].
+By integrating with the proposed decomposition stage, the adapted grasp pose directly modifies the variable block $\mathcal{B}_{i}$ into $\mathcal{B}^{\prime}_{i}$, enabling synthesis of new trajectories.
+IV-C Trajectory Modulation and Optimization
+After deconstructing the provided demonstration into blocks (Sec.
+IV-A) and adapting variable blocks via vision-based alignment (Sec.
+IV-B), the synthesized trajectory still requires two critical refinements to ensure executability and robustness as expalined below:
+IV-C1 Collision-Aware Validation
 Each adapted block undergoes kinematic feasibility checks.
-First, we validate the reachability of target poses $\mathbf{P}_{\textrm{grasp,novel}}$ using Inverse Kinematics (IK), ensuring the robot joint limits are satisfied.
+First, we validate the reachability of target poses $\mathbf{P}_{\textrm{\scriptsize grasp,novel}}$ using Inverse Kinematics (IK), ensuring the robot joint limits are satisfied.
 If the result returned is singular or unreachable (rarely happens), we will delete this sample.
-Second, a motion planner chitta2012moveit; schulman2014motion verifies potential collision between arms and scene objects by solving for collision-free paths between the start and end states in each block.
+Second, a motion planner [19, 75] verifies potential collision between arms and scene objects by solving for collision-free paths between the start and end states in each block.
 This simplified two-point constraint reduces computational overhead while preserving safety, assuming the trajectory of original given demonstration is collision-free.
-Instance-Level Motion Adaptation.
-To handle geometric variations across object instances with differing shapes and sizes, we adjust motion primitives in variable blocks based on the object 3D bounding box dimensions.
-Let $\Delta l$, $\Delta w$ and $\Delta h$ denote the length, width and height differences between the novel object bounding box $\textbf{b}_{\textrm{novel}}$ and the original $\textbf{b}_{\textrm{demo}}$.
+IV-C2 Instance-Level Motion Adaptation
+To handle geometric variations across object instances with differing shapes, we adjust motion primitives in variable blocks based on the object 3D bounding box dimensions.
+Let $\Delta l$, $\Delta w$ and $\Delta h$ denote the length, width and height differences between the novel object bounding box $\textbf{b}_{\textrm{\scriptsize novel}}$ and the original $\textbf{b}_{\textrm{\scriptsize demo}}$.
 Motion endpoints (e.g., grasp or release positions) are offset by:
@@
 For irregular shapes, depth-aware masks $M$ extracted in Sec.
-4.2 can approximate volumetric differences, or minimal human input provides precise measurements.
+IV-B can approximate volumetric differences, or minimal human input provides precise measurements.
 This hierarchical optimization stage operates at the block level rather than full trajectories, and enables efficient scaling while preserving the task semantics encoded in invariant blocks.
-Finally, validated and adapted blocks are recombined into synthetic trajectories $\tau^{\prime}\!=\!\{\tau_{\textrm{inv}},\{\mathcal{B}^{\prime}_{i}\}\}$, populating diverse and physically feasible demonstrations $\mathcal{D}_{syn}$.
-5 Experiments
-Our experiments address three core questions.
+Finally, validated and adapted blocks are recombined into synthetic trajectories $\tau^{\prime}\!=\!\{\tau_{\textrm{\scriptsize inv}},\{\mathcal{B}^{\prime}_{i}\}\}$, populating diverse and physically feasible demonstrations $\mathcal{D}_{syn}$.
+V Experiments
+Our experiments address five core questions.
 $\mathsf{Q}1$: Is BiDemoSyn truly efficient and user-friendly?
@@
 $\mathsf{Q}3$: Does synthetic demonstrations generalize to spatial and object variations?
-5.1 Experimental Setup and Protocol
-Tasks.
-We evaluate on six bimanual manipulation tasks (refer Fig.
-9) requiring contact-rich coordination: plugpen, inserting, unscrew, pouring, pressing and reorient.
-These tasks cover diverse primitive skills (e.g., grasping, rotating and handover), and involve both rigid and articulated objects.
-The one-shot demonstration for each task is collected via kinesthetic teaching.
-For the platform (refer Fig.
-8), two fixed-base arms equipped with parallel grippers are used.
-Scene perception is provided by a stereo camera capturing binocular images.
-More details of all tasks and the corresponding data collection process are in Appendix Sec.
-A.
-Policies & Baselines.
-For a comprehensive comparison, baselines contain two categories.
-One category is purely for data collection including point cloud editing (DemoGen xue2025demogen), real robot auto-rollout (YOTO zhou2025you), and human drag teaching (close to Teleoperation).
-Generally, the quality of collected demonstrations by these baselines is better in turn, but the cost is more time-consuming.
-After preparing the training data, we adapt three advanced visuomotor policies (DP chi2023diffusion, DP3 ze2024dp3 and EquiBot yang2024equibot) to bimanual settings by modifying their modeling spaces to dual-arm actions (refer Sec.
-3).
-Observation inputs are RGB-only images or segmented 3D point clouds of task-relevant objects.
-Final trained policies operate in the open-loop discrete keyposes prediction to align with our synthesized demonstration format.
-The another category is directly for bimanual manipulation without retraining including the zero-shot ReKep huang2024rekep, an advanced ReKep+ with oracle-level grasp labels at the beginning, and one-shot ODIL wang2025one.
-Metrics.
-We evaluate (1) synthesis efficiency: average time per demo over 50 tests, (2) data quality: visual authenticity of newly acquired observations, (3) success rates: real robot deployment with 30 trials per task, and (4) generalization: performance on unseen object placements and geometries.
+$\mathsf{Q}4$: Can BiDemoSyn seamlessly integrate with few-shot demonstrations to further improve robustness and data efficiency?
+$\mathsf{Q}5$: Do policies trained on BiDemoSyn data transfer zero-shot across different robot embodiments while maintaining high success rates?
 Figure 3: (A) The data collection efficiency comparison of different baselines and our BiDemoSyn, and (B) the generated data quality illustration of DemoGen.
 Although DemoGen has the highest synthesis efficiency, it cannot avoid visual artifacts caused by perspective transformation, so its data quality is the lowest.
-Our method can achieve a balance between the speed and quality.
-5.2 Comparison and Presentation of Results
-We here answer three questions raised earlier, demonstrating that BiDemoSyn can efficiently synthesize high-quality real-world demonstrations, enabling scalable and generalizable visuomotor policy training with minimal human input.
-Then there are some snapshots of real robot execution effects.
-($\mathsf{A}1$): The efficiency and usability of BiDemoSyn have obvious advantages over baselines.
-Without bells and whistles, our BiDemoSyn demonstrates superior efficiency and usability as illustrated in Fig.
-4.
-It requires about 5 seconds per demonstration to synthesize new trajectories, where primarily involving manual object repositioning and visual verification of initial frame alignment, followed by the trajectory optimization.
-In contrast, DemoGen generates trajectories in less then 1 second via scripted edits but suffers from perspective distortion artifacts, degrading data quality.
-The specific qualitative effect can be compared with Fig.
-4A and our results in Fig.
-2 and Fig.
-9.
-By relying on manual alignment and auto-replay, YOTO takes 42 seconds per demo which is equivalent to real robot execution time.
-While, teleoperation demands 91 seconds per demo for near-perfect but labor-intensive collection.
-These results highlight BiDemoSyn’s unique balance of speed (automated optimization) and reliability (artifact-free synthesis), making it the only method scalable to thousands of demonstrations without compromising real-world fidelity.
-Table 1: Quantitative comparison results of success rates for different methods under in-distribution and out-of-distribution evaluations. The # means the BiDemoSyn training data size.
-  in-distribution evaluations out-of-distribution evaluations
-Policy
-(Input
-Mode)
-Method
+Our method can achieve a good speed-quality balance.
+TABLE I: Quantitative results of different methods under in-distribution (ID) and out-of-distribution (OOD) evaluations.
+The OOD means removing one or a pair of objects from the training set for testing purposes only.
+The # means the training size.
+  in-distribution (ID) evaluations out-of-distribution (OOD) evaluations
+Policy (Input Mode) Method
 plugpen
@@
 reorient
-Average
-Success
-Rate
+Average Success Rate
 plugpen
@@
 reorient
-Average
-Success
-Rate
+Average Success Rate
 #3888 #7776 #1152 #2592 #1296 #1008 #2916 #3888 #1008 #972 #324 #756
-Training
--Free
-ReKep 15/30 13/30 12/30 15/30 12/30 09/30 42.2% 12/30 10/30 10/30 09/30 09/30 06/30 31.1%
+  Training -Free (RGB) ReKep 15/30 13/30 12/30 15/30 12/30 09/30 42.2% 12/30 10/30 10/30 09/30 09/30 06/30 31.1%
 ReKep+ 17/30 17/30 14/30 19/30 20/30 11/30 54.4% 15/30 14/30 12/30 12/30 11/30 09/30 40.6%
 ODIL 18/30 19/30 13/30 18/30 18/30 12/30 54.4% 12/30 14/30 08/30 11/30 13/30 08/30 36.7%
-DP
-DemoGen 18/30 20/30 16/30 16/30 14/30 15/30 55.0% 08/30 03/30 10/30 04/30 03/30 07/30 19.4%
+MAGIC 19/30 18/30 14/30 18/30 19/30 12/30 55.6% 13/30 14/30 10/30 12/30 13/30 09/30 39.4%
+DP (RGB) DemoGen 18/30 20/30 16/30 16/30 14/30 15/30 55.0% 08/30 03/30 10/30 04/30 03/30 07/30 19.4%
 YOTO 19/30 20/30 16/30 17/30 14/30 16/30 56.7% 08/30 04/30 10/30 05/30 04/30 07/30 21.1%
 BiDemoSyn 22/30 24/30 22/30 21/30 17/30 19/30 67.8% 13/30 10/30 18/30 15/30 08/30 12/30 42.2%
-DP3
-DemoGen 21/30 24/30 19/30 20/30 18/30 17/30 66.1% 11/30 06/30 14/30 07/30 05/30 11/30 30.0%
+DP3 (PCD) DemoGen 21/30 24/30 19/30 20/30 18/30 17/30 66.1% 11/30 06/30 14/30 07/30 05/30 11/30 30.0%
 YOTO 22/30 23/30 20/30 20/30 18/30 19/30 67.8% 12/30 06/30 15/30 08/30 06/30 14/30 33.9%
 BiDemoSyn 26/30 28/30 25/30 24/30 21/30 22/30 81.1% 16/30 14/30 21/30 18/30 11/30 18/30 54.4%
-EquiBot
-DemoGen 22/30 24/30 20/30 20/30 19/30 19/30 68.9% 13/30 10/30 16/30 11/30 10/30 13/30 40.6%
+EquiBot (PCD) DemoGen 22/30 24/30 20/30 20/30 19/30 19/30 68.9% 13/30 10/30 16/30 11/30 10/30 13/30 40.6%
 YOTO 23/30 24/30 20/30 21/30 20/30 19/30 71.1% 15/30 11/30 18/30 14/30 12/30 16/30 47.8%
 BiDemoSyn 28/30 28/30 26/30 25/30 24/30 25/30 86.7% 18/30 20/30 24/30 21/30 17/30 20/30 66.7%
-Figure 5: Comparison between training data scale and the success rate. Less sized data is randomly sampled out of the total dataset at the task level. DP3 is chosen as the visuomotor policy.
+V-A Experimental Setup and Protocol
+Tasks.
+We evaluate on six bimanual manipulation tasks requiring contact-rich coordination: plugpen, inserting, unscrew, pouring, pressing and reorient.
+These tasks cover diverse primitive skills (e.g., grasping, rotating and handover), and involve both rigid and articulated objects.
+The one-shot demonstration for each task is collected via kinesthetic teaching.
+For the platform, two fixed-base arms equipped with parallel grippers are used as the main workspace.
+An auxiliary humanoid dual-arm platform was also used to verify the cross-platform deployment capability of the learned policy.
+Scene perception is provided by a stereo camera capturing binocular images.
+More details of tasks, harwares and data collection are in Supplementary Materials.
+Policies and Baselines.
+Compared baselines contain two categories.
+One category is purely for data collection including point cloud editing (DemoGen [93]), real robot auto-rollout (YOTO [107]), and human drag teaching (close to Teleoperation).
+Generally, the quality of collected demonstrations by these baselines is better in turn, but the cost is more time-consuming.
+After preparing the training data, we adapt three advanced visuomotor policies (DP [16], DP3 [98] and EquiBot [94]) to bimanual settings by modifying their modeling spaces to dual-arm actions (refer Sec.
+III).
+Observation inputs are RGB-only images or segmented 3D point clouds of task-relevant objects.
+Final trained policies operate in the open-loop discrete keyposes prediction to align with our synthesized demonstration format.
+The another category is directly for bimanual manipulation without retraining including the zero-shot ReKep [38], an advanced ReKep+ with oracle-level grasp labels at the beginning, and two one-shot imitation learning methods ODIL [84] and MAGIC [60].
+Metrics.
+We evaluate (1) synthesis efficiency: average time per demo over 50 tests, (2) data quality: visual authenticity of newly acquired observations, (3) success rates: real robot deployment with 30 trials per task, and (4) generalization: performance on unseen object placements and geometries.
+V-B Comparison and Presentation of Results
+We here answer three questions raised earlier, demonstrating that BiDemoSyn can efficiently synthesize high-quality real-world demonstrations, enabling scalable and generalizable visuomotor policy training with minimal human input.
+Then there are some snapshots of real robot execution effects.
+Figure 4: Comparison between training scale and success rate. Less sized data is randomly sampled out of the total dataset at task-level. DP3 is chosen as the visuomotor policy.
+Figure 5: Analysis of spatial generalization for variations in (A) position sampling density and (B) orientation sampling diversity. DP3 is chosen as the policy.
+Figure 6: Visualization of all six bimanual tasks performed on real robots.
+All models are trained and tested under the ID evaluations.
+EquiBot is chosen as the visuomotor policy.
+Key highlighted dual-arm coordination movements associated with each task are partially enlarged at the bottom row for quick review.
+($\mathsf{A}1$): The efficiency and usability of BiDemoSyn have obvious advantages over baselines.
+Without bells and whistles, our BiDemoSyn demonstrates superior efficiency and usability as illustrated in Fig.
+3.
+It requires about 5 seconds per demonstration to synthesize new trajectories, where primarily involving manual object repositioning and visual verification of initial frame alignment, followed by the trajectory optimization.
+In contrast, DemoGen generates trajectories in less then 1 second via scripted edits but suffers from perspective distortion artifacts, degrading data quality.
+The specific qualitative effect can be compared with Fig.
+3A and our results in Fig.
+1.
+By relying on manual alignment and auto-replay, YOTO takes 42 seconds per demo which is equivalent to real robot execution time.
+While, teleoperation demands 91 seconds per demo for near-perfect but labor-intensive collection.
+These results highlight BiDemoSyn’s unique balance of speed (automated optimization) and reliability (artifact-free synthesis), making it the only method scalable to thousands of demonstrations without compromising real-world fidelity.
 ($\mathsf{A}2$): Demonstrations obtained via BiDemoSyn can support scalable imitation learning.
-We applied BiDemoSyn for efficient collection of thousands demonstrations per task (Fig.
-9 right), densely covering the workspace with instance-level object diversity, thereby providing abundant training data for imitation learning.
-For two reproduced baselines, DemoGen xue2025demogen matches our data scale but suffers from visual artifacts, while YOTO zhou2025you collects only 1/10 the data per task for being limited by its time-consuming replay mechanism.
-After obtaining adequate data, we trained three advanced visuomotor policies DP chi2023diffusion, DP3 ze2024dp3 and EquiBot yang2024equibot.
+We applied BiDemoSyn for efficient collection of thousands demonstrations per task, densely covering the workspace with instance-level object diversity, thereby providing abundant training data for imitation learning.
+For two reproduced baselines, DemoGen [93] matches our data scale but suffers from visual artifacts, while YOTO [107] collects only 1/10 the data per task for being limited by its time-consuming replay mechanism.
+After obtaining adequate data, we trained three advanced visuomotor policies DP [16], DP3 [98] and EquiBot [94].
 As shown in Tab.
-1 left, BiDemoSyn achieves superior in-distribution (ID) success rates across all six tasks, outperforming both baselines regardless of the policy architecture.
+I left, BiDemoSyn achieves superior in-distribution (ID) success rates across all six tasks, outperforming both baselines regardless of the policy architecture.
 This confirms that BiDemoSyn’s artifact-free synthesis and efficient data scaling reliably support the scaling laws of imitation learning.
-Besides, it is unsurprising that policies trained with demonstrations consistently outperform the two training-free baselines ReKephuang2024rekep and ODIL wang2025one.
-To further quantify scalability, we evaluated training data size-performance relationships on all six tasks of all methods.
-As shown in Fig.
+Besides, it is unsurprising that policies trained with demonstrations consistently outperform three training-free baselines ReKep[38], ODIL [84], and MAGIC [60].
+To further quantify scalability, we evaluated training size-performance relationships on six tasks of all methods.
+As in Fig.
 5, BiDemoSyn-trained policies exhibit stronger scaling trends than DemoGen (e.g., 60% vs.
-39% at 1106 demos), aligning with community observations that data quality and quantity jointly drive visuomotor policy performance liu2025rdt; black2024pi0; pertsch2025fast; lin2025data.
-Figure 6: Analysis of spatial generalization, including variations in (A) position sampling density and (B) orientation sampling diversity. DP3 is chosen as the visuomotor policy.
+39% at 1106 demos), aligning with community observations that data quality and quantity jointly drive visuomotor policy performance [58, 5, 71, 53].
+TABLE II: Quantitative results for few-shot BiDemoSyn.
+Although the number of training demonstrations is the same for each setting, the number of seed demonstrations for trajectory synthesis is different.
+The # means the training size.
+  in-distribution (ID) evaluations out-of-distribution (OOD) evaluations
+Trained Policy Few-Shot BiDemoSyn
+plugpen
+inserting
+unscrew
+pouring
+pressing
+reorient
+Average Success Rate
+plugpen
+inserting
+unscrew
+pouring
+pressing
+reorient
+Average Success Rate
+#3888 #7776 #1152 #2592 #1296 #1008 #2916 #3888 #1008 #972 #324 #756
+  EquiBot 1-shot 28/30 28/30 26/30 25/30 24/30 25/30 86.7% 18/30 20/30 24/30 21/30 17/30 20/30 66.7%
+5-shot 28/30 28/30 26/30 26/30 25/30 25/30 87.8% 20/30 21/30 24/30 23/30 18/30 21/30 70.0%
+10-shot 28/30 28/30 26/30 27/30 25/30 26/30 88.9% 21/30 21/30 25/30 23/30 18/30 22/30 71.7%
+15-shot 28/30 28/30 27/30 27/30 26/30 26/30 90.0% 21/30 22/30 25/30 23/30 18/30 22/30 72.2%
+20-shot 28/30 28/30 27/30 27/30 26/30 26/30 90.0% 21/30 22/30 25/30 23/30 18/30 22/30 72.2%
 ($\mathsf{A}3$): Policies trained on BiDemoSyn data can achieve better generalization to unseen variations.
@@
 As shown in Tab.
-1 right, policies trained on BiDemoSyn data consistently achieve higher success rates compared to all baselines in out-of-distribution (OOD) settings.
+I right, policies trained on BiDemoSyn data consistently achieve higher success rates compared to all baselines in out-of-distribution (OOD) settings.
 While all methods exhibit performance drops relative to ID evaluations, BiDemoSyn’s superior OOD robustness highlights its ability to capture task-invariant features through vision-aligned synthesis.
-To further analyze spatial generalization, we vary the density of positional and orientation coverage in training data: for unscrew, we synthesize trajectories with sparse-to-dense workspace coverage (Fig.
-6A), and for reorient, we incrementally increase the diversity of object orientations (Fig.
-6B).
+To further analyze spatial generalization, we vary the density of positional and orientation coverage in training data: for unscrew, we synthesize trajectories with sparse-to-dense workspace coverage (see Fig.
+5A), and for reorient, we incrementally increase the diversity of object orientations (see Fig.
+5B).
 To exclude the effect of data size differences, the total number of train-set is always augmented to a standard size (e.g., #1152 and #1008) via the DemoGen to edit some observations at small distances.
@@
 These results confirm that BiDemoSyn’s ability to cheaply synthesize positionally and orientationally diverse demos directly translates to enhanced policy generalization, aligning with empirical insights that broad data coverage mitigates spatial distribution shifts.
-Figure 7: Visualization of all six bimanual tasks performed on real robots.
-All models are trained and tested under the ID evaluations.
-EquiBot is chosen as the visuomotor policy.
-Key dual-arm coordination movements associated with each task are partially enlarged for quick review.
-Qualitative Results: Real rollouts in Fig.
-7 showcase effectiveness across six tasks.
+Qualitative Results: Fig.
+6 shows effectiveness across six tasks.
 For example, the plugpen policy precisely aligns and inserts the pen into the narrow cap even with $\pm$5mm positional noise.
@@
 For pressing, predicted actions reliably activate nozzles without over-pressuring, while reorient handles object handovers and flips with smooth dual-arm transitions.
-These visualizations corroborate our quantitative findings, emphasizing BiDemoSyn’s ability to handle real-world complexity in contact-rich manipulation.
-6 Conclusion and Limitation
-We present BiDemoSyn, a framework for synthesizing real-world bimanual demonstrations from a single human example.
-By decomposing tasks into invariant and adaptable components, leveraging vision-guided scene alignment, and hierarchically optimizing trajectories, BiDemoSyn generates diverse, collision-free data without synthetic gaps.
-Experiments on six complex contact-rich tasks demonstrate its efficiency, scalability, and generalization, outperforming strong baselines and pure teleoperation.
+These results corroborate our quantitative findings, emphasizing BiDemoSyn’s ability to handle real-world complexity in contact-rich cases.
+Figure 7: Visualization for the cross-embodiment deployment of BiDemoSyn-trained policies on an unseen humanoid-style dual-arm robot for the plugpen and reorient tasks.
+($\mathsf{A}4$): BiDemoSyn can seamlessly integrate with few-shot demonstrations to improve robustness.
+Although BiDemoSyn is formulated in a one-shot paradigm, it can seamlessly incorporate additional real demonstrations (5–20 samples) as complementary seeds without altering the synthesis pipeline.
+We summarized this ablation results in Tab.
+II.
+Empirically, we can find that few-shot extensions provide limited gains under ID settings, where one-shot–generated datasets already yield near-saturated performance on seen settings.
+In contrast, moderate but consistent improvements are observed under OOD conditions, primarily due to expanded coverage of object geometries and intermediate contact states introduced by new demonstrations.
+These results suggest that BiDemoSyn’s one-shot synthesis already produces high-quality, physically grounded demonstrations, while few-shot seeds mainly contribute by enriching representational diversity rather than correcting systematic failure modes.
+Overall, this study confirms that BiDemoSyn offers a favorable trade-off: it remains highly effective in the one-shot regime, benefits selectively from few-shot data in OOD scenarios, and supports hybrid one-shot + few-shot usage without increasing complexity.
+($\mathsf{A}5$): Policies trained on BiDemoSyn data transfer effectively across robot embodiments in a zero-shot manner.
+BiDemoSyn facilitates zero-shot cross-embodiment policy deployment by design, owing to its object-centric observation formulation and the use of end-effector–centric 6-DoF action representation.
+These choices decouple policy learning from robot-specific kinematics and enable direct transfer to new hardware platforms.
+To evaluate this capability, we deploy visuomotor policies trained on the primary dual-arm platform to a previously unseen humanoid-style bimanual robot, without additional training or fine-tuning.
+We consider two representative tasks, plugpen and reorient, using EquiBot policies trained under ID settings.
+As shown in Fig.
+7, the transferred policies successfully reproduce precise bimanual alignment for insertion and reorientation.
+Quantitative evaluations indicate that success rates on the new embodiment are comparable to those reported in Tab.
+I left.
+These results demonstrate that BiDemoSyn-trained policies exhibit strong embodiment-agnostic behavior, highlighting the potential for scalable deployment across heterogeneous robotic platforms.
+VI Conclusion and Limitation
+We introduce BiDemoSyn, a real-world demonstration synthesis framework that enables scalable bimanual manipulation learning from a single kinesthetic example.
+By decomposing demonstrations into invariant coordination blocks and object-dependent adaptations, and instantiating them via vision-based object anchoring and trajectory optimization, BiDemoSyn generates large-scale, physically feasible demonstrations without simulation or repeated teleoperation.
+Across diverse bimanual tasks, policies trained on BiDemoSyn data achieve strong generalization to novel objects, long-horizon compositions, and unseen robot embodiments.
+We further show that the framework naturally extends to few-shot synthesis, improving robustness with minimal additional supervision.
 This work establishes a new paradigm for scalable imitation learning, bridging the divide between data quantity and physical fidelity in bimanual robotic manipulation.
-Limitations: While BiDemoSyn excels in static, geometrically varied scenes, it faces challenges in dynamic environments and extreme shape variations (e.g., deformable fabric objects).
+Limitations: While BiDemoSyn excels in static, geometrically varied scenes, it faces challenges in fully dynamic environments and extreme shape variations (e.g., highly deformable or articulated objects).
 The current hierarchical optimization assumes quasi-static motions, limiting its applicability to highly dynamic tasks like catching.
@@
 Appendix
-This supplementary part provides detailed explanations and extensions to the main paper.
+This supplementary provides detailed explanations and extensions to the main paper.
 Sec.
-A outlines the design rationale for our six bimanual manipulation tasks, including platform specifications, task-specific objectives, and one-shot demonstration acquisition protocols.
+VII outlines the design rationale of six bimanual manipulation tasks, including platform specifications, task-specific objectives, and one-shot demonstration acquisition protocols.
 Sec.
-B details the data collection standards, UI-assisted collection tools, and post-processing workflows.
+VIII details the data collection standards, UI-assisted collection tools, and post-processing workflows.
 These constitute the trajectory synthesis pipelines for generating real-world demonstrations with BiDemoSyn.
-We have also added more details about BiDemoSyn to help readers better understand this study.
+We have also added more details about BiDemoSyn to help to understand it.
 Sec.
-C elaborates on policy training and deployment, covering implementation specifics for baseline methods (DemoGen xue2025demogen, YOTO zhou2025you) and diffusion policies (DP3 ze2024dp3, EquiBot yang2024equibot), alongside real-robot deployment.
+IX elaborates on policy training, covering implementation specifics for baselines (DemoGen [93], YOTO [107]) and diffusion policies (DP3 [98], EquiBot [94]), alongside real-robot deployment.
 Sec.
-D provides additional visualizations of real-robot executions and failure case analyses, offering insights into current limitations.
+X provides additional visualizations of real-robot executions and failure case analyses, offering insights into current limitations.
 Sec.
-E concludes with reflections and future directions, aiming to extend BiDemoSyn to challenge more complex tasks (e.g., deformable or dynamic manipulation) and enhance cross-domain generalization.
-Sec.
-F is the statement of the use of Large Language Models.
-Appendix A Design of Bimanual Manipulation Tasks
-Figure 8: Left: The fixed-base dual-arm manipulator platform (a table with two robot arms, two grippers and the binocular camera) used in this research.
-Right: The object assets involved in our six bimanual manipulation tasks.
-All objects have been scaled down proportionally.
-A.1 Hardware and Platform
+XI concludes with reflections and future directions, aiming to extend BiDemoSyn to challenge more complex tasks and enhance cross-domain generalization.
+VII Design of Bimanual Manipulation Tasks
+VII-A Hardware and Platform
 […]
-A.2 Bimanual Task Formulation
-We design six bimanual manipulation tasks (including plugpen, inserting, unscrew, pouring, pressing and reorient) to comprehensively evaluate dual-arm coordination and generalization.
-Each task involves at least two category-level object instances (refer Fig.
-8 right) and requires diverse primitive skills, such as single-arm actions (grasping, placing, carrying, precision rotation) and dual-arm coordination (fixed-twisting, plug-in, handover, alignment).
-All tasks are long-horizon, starting from goal-conditioned initial grasping (objects fully separated from robots), contrasting prior works already grasping/holding the manipulated object and focusing on short-horizon atomic skills such as untwisting lin2024twisting; bahety2024screwmimic or handover huang2023dynamic.
+Figure 8: (Top) Primary.
+The fixed-base dual-arm platform having a table with two robot arms, two grippers and the binocular camera.
+(Bottom) Auxiliary.
+The humanoid dual-arm platform for validating cross-embodiment capabilities.
+VII-B Bimanual Task Formulation
+We design six bimanual tasks (plugpen, inserting, unscrew, pouring, pressing and reorient) to comprehensively evaluate dual-arm coordination and generalization.
+Each task involves at least two category-level object instances (Fig.
+9) and requires diverse primitive skills, such as single-arm actions (grasping, placing, precision rotation) and dual-arm coordination (fixed-twisting, plug-in, handover).
+All tasks are inherently long-horizon, starting from goal-conditioned initial grasping (objects fully separated from robots), contrasting prior works already grasping/holding the manipulated object and focusing on short-horizon atomic skills such as untwisting [54, 1] or handover [37].
 Fig.
@@
 Below we detail each task:
-[…]
+Figure 9: Object assets involved in six bimanual manipulation tasks. All objects have been scaled down proportionally.
   • plugpen: plug the pen with the pen cap.
@@
 Steps: Right arm vertically presses the nozzle; left arm tilts the cup to catch water.
-Challenges: Nozzle positioning accuracy ($\pm$5mm), open-loop force control.
+Challenges: Nozzle positioning accuracy ($\pm$5mm), open-loop fine-grained force control.
   • reorient: flip the spoon so that the bottom is facing up.
@@
 Challenges: Unstable grasp during handover, rotational precision for flipping ($\pm$10).
+[…]
 These tasks collectively stress spatial reasoning, contact-aware coordination, and generalization to instance-level variations, which are cornerstones of real-world bimanual manipulation.
@@
 (3) Third-person view occlusions occasionally hinder precise alignment (e.g., cannot discern the groove side of the pen cap).
-A.3 Obtaining and Decomposing the One-Shot Demonstration
+VII-C Obtaining and Decomposing the One-Shot Demonstration
 As stated in the main content, we collect one-shot demonstrations via kinesthetic teaching: an operator manually guides both arms through task-critical waypoints, recording the 6DoF end-effector poses (relative to each robot base frame) and gripper binary states at each pause.
@@
 During kinesthetic teaching, the demonstrator physically guides the two manipulators through the task while pausing at intuitive semantic boundaries, such as after a stable grasp, after completing an alignment adjustment, or before switching from transportation to interaction.
-These pauses, which arise organically from human motion patterns, produce clear temporal discontinuities that align well with the block-based representation introduced in Sec.
-4.1 of the main text.
+These pauses, which arise organically from human motion patterns, produce clear temporal discontinuities that align well with the block-based representation introduced in the main paper.
 Importantly, this process does not require any manual annotation or labeling.
@@
 This makes the procedure no more labor-intensive than collecting a single ordinary demonstration.
-We explored alternative strategies for fully automatic segmentation, such as keyframe/keypose detection and curvature-based change-point identification james2022coarse; shridhar2023perceiver; ma2024hierarchical.
+We explored alternative strategies for fully automatic segmentation, such as keyframe/keypose detection and curvature-based change-point identification [39, 78, 61].
 However, reliably identifying semantically meaningful waypoints from a single demonstration proved brittle, especially when waypoints must correspond to subtle but crucial phases of contact establishment, re-grasping, or dual-arm synchronization.
-Hardware-assisted approaches (e.g., instrumented gloves as in DexCap wang2024dexcap or end-effector teaching interfaces as in UMI chi2024universal and DexUMI xu2025dexumi) can offer additional cues for segmentation but introduce alignment overhead, cross-morphology calibration challenges, and potential loss of semantic correspondence between human motion and robot embodiment.
+Hardware-assisted approaches (e.g., instrumented gloves as in DexCap [82] or end-effector teaching interfaces as in UMI [17] and DexUMI [92]) can offer additional cues for segmentation but introduce alignment overhead, cross-morphology calibration challenges, and potential loss of semantic correspondence between human motion and robot embodiment.
 Given these trade-offs, the kinesthetic-teaching-driven segmentation remains a pragmatic and robust choice: it imposes no additional annotation burden, preserves semantic alignment across blocks, and provides boundaries well suited for constructing invariant and variant components for real-world synthesis.
-Appendix B Details of Data Collection, Processing and Synthesis
-Figure 10: The specific grid cell division way for each task. For tasks involving two manipulated objects, the total number of grid cells will be divided equally between the left side and right side.
-Figure 11: UI interface examples.
-When collecting data, the six tasks need to display different grid cells in real time, as well as perceive and track objects related to the tasks.
+VIII Details of Data Collection and Process
+Figure 11: The grid cell division for each task. For tasks involving two manipulated objects, the total number of grid cells will be divided equally between the left and right side.
+VIII-A Task-Oriented Collection Standards
+[…]
+Figure 12: UI interfaces.
+When collecting data, six tasks need to display different grid cells in real time, as well as perceive and track task-related objects.
 Note that these points, crosses and lines are drawn digitally, which are not marks in the real world.
-B.1 Task-Oriented Collection Standards
-[…]
-B.2 Collection UI Interface Development
+VIII-B Collection UI Interface Development
 To streamline data collection, we develop a real-time UI interface (Fig.
-11) that visualizes task-specific grid cells, dynamically tracks object positions/poses, and validates alignment with the one-shot demonstration.
-The interface overlays grid layouts onto live camera feeds, highlights detected objects (using YOLOE wang2025yoloe for fast detection and segmentation), and marks completed cells using red crosses.
+12) that visualizes task-specific grid cells, dynamically tracks object positions/poses, and validates alignment with the one-shot demonstration.
+The interface overlays grid layouts onto live camera feeds, highlights detected objects (using YOLOE [81] for fast detection and segmentation), and marks completed cells using red crosses.
 The position of objects is represented by red dots (some objects also use yellow arrows to indicate their orientation).
@@
 A sidebar displays positional/orientational deviations from the reference demonstration, enabling quick verification.
-If object perception fails, the system falls back to Florence2 xiao2024florence and SAM2 ravi2024sam for robust detection and segmentation.
-Operators trigger a ”capture-align-validate-save” cycle every 2 seconds.
+If object perception fails, the system falls back to Florence2 [89] and SAM2 [74] for robust detection and segmentation.
+Operators trigger a capture-align-validate-save cycle every 2 seconds.
 And successful saves prompt visual feedback (UI flash of corresponding grid cells), signaling readiness for the next sample.
@@
 Please refer our Supplemental Videos that demonstrate the vivid workflow, illustrating how the UI balances efficiency with precision.
-B.3 Post-Processing of Collected Data
-This part includes the last two stages of BiDemoSyn: initial frame alignment and trajectory optimization (defined in the method part of main paper).
+VIII-C Post-Processing of Collected Data
+This part includes the last two stages of BiDemoSyn: initial frame alignment and trajectory optimization.
 First, pre-recorded positional/orientational deviations (validated via the UI interface during collection) align novel observations with the one-shot demonstration reference frame using the rigid transformation.
@@
 Finally, validated trajectories are stored in $\mathcal{D}_{syn}$ to complete the synthesis pipeline.
-B.4 Clarification on Hyperparameters and Predefined Symbols
-In Sec.
-4, several task-specific thresholds and predefined parameters are introduced for clarity and reproducibility.
+VIII-D Hyperparameters and Predefined Symbols
+In the method section, several thresholds and parameters are introduced for clarity and reproducibility.
 Here we further explain their meaning and selection strategy.
@@
 The symbol $\zeta$ is a static tolerance threshold to segment motion boundaries under quasi-static assumptions.
-The symbol $\gamma$ is a refinement threshold used to create fine-grained AEPs after bounding block duration ($T_{\textrm{max}}$).
+The symbol $\gamma$ is a refinement threshold used to create fine-grained AEPs after bounding block duration ($T_{\textrm{\scriptsize max}}$).
 And the scaling factor $\lambda$ in Eqn.
@@
 Larger or smaller values tend to shift grasp points away from functional regions (e.g., bottle mid-section in unscrew), while $\lambda=1.0$ already suffices for most tasks.
-Similarly, thresholds for collision-checking or pose-alignment rejection (e.g., rejecting infeasible IK solutions or near-singular configurations) were determined through small-scale grid search and validated in pilot runs.
+Similarly, thresholds for collision-checking or pose-alignment rejection (e.g., rejecting infeasible IK solutions or near-singular samples) were determined through small-scale grid search and validated in pilot runs.
 Importantly, such hyperparameters are not per-trial tunable knobs, but rather fixed system constants, ensuring consistent data synthesis across tasks.
-B.5 Additional Evaluation of Intermediate Processes
+VIII-E Evaluation of Intermediate Processes
 We provide further evaluation to justify intermediate design choices from three aspects: Segmentation Quality, Pose Estimation Robustness and Impact on Data Reliability.
-(1) Firstly, object masks were obtained from Florence2 xiao2024florence and SAM2 ravi2024sam, two state-of-the-art vision foundation models (VFMs).
+(1) Firstly, object masks were obtained from Florence2 [89] and SAM2 [74], two state-of-the-art vision foundation models (VFMs).
 On held-out captures without foreground occlusion, the segmentation accuracy exceeded 97%, ensuring reliable geometric inputs for subsequent processing.
@@
 (3) Thirdly, in Sec.
-D.2 (Statistics and Analysis of Failed Cases), we showed that perception and orientation errors together account for  47% of failures, highlighting their importance.
-Nevertheless, these intermediate modules are already close to optimal under current open-loop constraints, which explains the high success rate ( 98%) of the data collection system.
-B.6 Positioning with Respect to Prior Work
-To avoid potential confusion with prior works such as ReKep huang2024rekep and ODIL wang2025one, we emphasize both the similarities and the fundamental differences.
+X-D, we showed that perception and orientation errors together account for $\sim$47% of failures, highlighting their importance.
+Nevertheless, these intermediate modules are already close to optimal under current open-loop constraints, which explains the high success rate ($\sim$98%) of the data collection system.
+VIII-F Positioning with Respect to Prior Works
+To avoid potential confusion with prior related works such as ReKep [38], ODIL [84] and MAGIC [60], we emphasize both the similarities and the fundamental differences.
 Like these methods, BiDemoSyn exploits one-shot demonstrations and compositional reasoning; however, its core objective is not direct policy execution, but rather rapid and scalable synthesis of real-world training demonstrations.
-This distinction is crucial: while prior methods target zero-/few-shot policy generalization (often trading reliability for flexibility), BiDemoSyn deliberately focuses on maximizing data quality and efficiency, enabling downstream training of robust visuomotor policies (e.g., DP3 ze2024dp3 and EquiBot yang2024equibot).
-Our evaluation against DemoGen xue2025demogen and YOTO zhou2025you, which are two strong baselines addressing the same bottleneck, demonstrates that BiDemoSyn achieves a superior balance between efficiency and data reliability, which is the practical need for scaling imitation learning.
-A concurrent work MoMaGen li2025momagen tackles scalable bimanual mobile manipulation by generating diverse trajectories in simulation, focusing on base reachability and camera visibility through constrained optimization.
+This distinction is crucial: while prior methods target zero-/few-shot policy generalization (often trading reliability for flexibility), BiDemoSyn deliberately focuses on maximizing data quality and efficiency, enabling downstream training of robust visuomotor policies (e.g., DP3 [98] and EquiBot [94]).
+Our evaluation against DemoGen [93] and YOTO [107], which are two strong baselines addressing the same bottleneck, demonstrates that BiDemoSyn achieves a superior balance between efficiency and data reliability, which is the practical need for scaling imitation learning.
+A concurrent work MoMaGen [49] tackles scalable bimanual mobile manipulation by generating diverse trajectories in simulation, focusing on base reachability and camera visibility through constrained optimization.
 Although sharing a high-level goal of reducing human effort, MoMaGen and BiDemoSyn operate under fundamentally different assumptions: MoMaGen relies on noise-free simulation assets and virtual scene perturbations, whereas BiDemoSyn performs real-world demonstration synthesis, directly addressing physical-domain challenges such as noisy perception, contact feasibility, and one-shot decomposition under real sensor observations.
@@
 Thus, BiDemoSyn should be viewed not as a direct competitor to simulation-based frameworks like MoMaGen, but as a complementary real-world solution specifically designed for scalable, high-fidelity data generation where simulation pipelines cannot fully capture real-world physical constraints.
-B.7 Evaluation of Demonstration Quality
+VIII-G Evaluation of Demonstration Quality
 Evaluating the quality of synthesized demonstrations is essential because the downstream visuomotor policies depend on both the fidelity of observations and the physical feasibility of actions.
@@
 Teleoperation and BiDemoSyn preserve raw RGB-D observations, ensuring perfect consistency with the physical environment.
-YOTO zhou2025you and DemoGen xue2025demogen modify the seed point cloud to reposition objects, which introduces geometric distortions and depth inconsistencies, particularly as the displacement grows (see Fig.
-4B).
+YOTO [107] and DemoGen [93] modify the seed point cloud to reposition objects, which introduces geometric distortions and depth inconsistencies, particularly as the displacement grows (see Fig.
+4B in the main paper).
 These artifacts degrade the policy’s perception module by coupling unrealistic geometry with real sensor noise.
-Trajectory Feasibility and Physical Grounding.
+Trajectory Feasibility & Physical Grounding.
 A demonstration is considered high-quality if the associated action sequence: (1) yields a collision-free, dynamically stable trajectory, (2) respects the contact geometry implied by the task semantics, and (3) can be executed on the real robot without additional correction or replanning.
-Teleoperation and YOTO zhou2025you provide ground-truth trajectories.
+Teleoperation and YOTO [107] provide ground-truth trajectories.
 BiDemoSyn produces partially reused (invariant blocks) and partially adapted (variant blocks) trajectories aligned with real-world object pose and geometry.
-DemoGen xue2025demogen synthesizes trajectories directly from edited scene representations, which can deviate from physically stable contact geometries.
+DemoGen [93] synthesizes trajectories directly from edited scene representations, which can deviate from physically stable contact geometries.
 Finally, Fig.
-4A reflects the interplay of these two components.
-Demonstrations with high-quality RGB-D observations and physically feasible actions rank highest.
+4A in the main paper reflects the interplay of these two components.
+Demonstrations with high-quality observations and physically feasible actions rank highest.
 This ordering also predicts downstream policy performance: policies trained with cleaner, physically grounded demonstrations demonstrate higher success rates and better robustness under OOD evaluation.
 These expanded analyses clarify the qualitative and quantitative factors driving the reported comparisons and support the role of demonstration quality as a key determinant of real-world imitation learning performance.
-B.8 Diversity of Synthetic Demonstrations
+VIII-H Diversity of Synthetic Demonstrations
 A natural question is whether additional diversity (such as reordering or randomizing block sequences) would improve the effectiveness of the synthesized dataset.
 In BiDemoSyn, we intentionally do not introduce such sequence-level variations.
-Our empirical findings align with recent analyses in large-scale teleoperation datasets focusing on the diversity exploration and ablation for scalable robotic manipulation shi2025diversity, which show that low-level behavioral variability (e.g., minor differences in grasp location, end-effector speed profiles, or redundant micro-motions) does not necessarily benefit policy learning and may even hinder convergence.
-Instead, we adopt the principle that task success is primarily determined by achieving the correct functional interactions, not by how those interactions are executed.
-Accordingly, in the design of BiDemoSyn, we maintain the original causal structure of the task by preserving the block ordering extracted from the one-shot demonstration, while introducing diversity along the dimensions that matter for generalization, namely: (1) spatial variation of object pose (translation + rotation), and (2) category-level instance variation (shape and size differences within the same object class).
-These forms of variation are directly tied to visuomotor generalization and produce clear benefits in downstream policy performance (see Tab.
-1).
+Our empirical findings align with recent analyses in large-scale teleoperation datasets focusing on the diversity exploration and ablation for scalable robotic manipulation [77], which show that low-level behavioral variability (e.g., minor differences in grasp location, end-effector speed profiles, or redundant micro-motions) does not necessarily benefit policy learning and may even hinder convergence.
+Instead, we adopt the principle that task success is primarily determined by achieving correct interactions, not by how those interactions are executed.
+Accordingly, in BiDemoSyn’s design, we maintain the original causal structure of the task by preserving the block ordering extracted from the one-shot demonstration, while introducing diversity along the dimensions that matter for generalization, namely: (1) spatial variation of object pose (translation + rotation), and (2) category-level instance variation (shape and size differences within the same object class).
+These forms of variation are directly tied to visuomotor generalization and produce clear benefits in downstream policy performance.
 In contrast, randomizing the internal block structure introduces behavioral diversity that is orthogonal to task semantics and does not improve success rates in our preliminary tests.
 We therefore avoid unnecessary variation and instead focus on environment-induced diversity, which is both meaningful and consistent with the underlying task geometry.
-Appendix C Training and Deployment of Imitation Policies
-C.1 Reproduction of Baselines
+Figure 13: The adapted visuomotor network architecture for imitation learning of bimanual tasks.
+The input visual observation is simplified to a point cloud of manipulated objects.
+Note that the vision encoder is slightly different in DP3 [98] and EquiBot [94].
+The illustrated one is for EquiBot.
+Figure 14: Qualitative real robot rollout samples for defined six bimanual tasks. Best to view after zooming in.
+Figure 15: Qualitative real robot rollout examples for two newly added long-horizon tasks (pouring and reorient+unscrew).
+Figure 16: Qualitative examples of applying dynamic interferences during the pre-grasping execution.
+From top to bottom, they are segments of the dynamic closed-loop grasping phase of tasks pouring and reorient+unscrew, where each object is manually disturbed from one to three times.
+The red arrow indicates the direction of the manually moved object (interfering).
+The cyan arrow and yellow arrow indicate the movement direction of the left and right robotic arms (chasing), respectively.
+IX Training and Deployment of Imitation Policies
+IX-A Reproduction of Baselines
 […]
-Figure 12: The adapted visuomotor policy network architecture for imitation learning of bimanual manipulation tasks.
-The input visual observation is simplified to a point cloud of manipulated objects.
-Note that the vision encoder is slightly different in DP3 ze2024dp3 and EquiBot yang2024equibot.
-The illustrated one is for EquiBot.
-C.2 Implementation of Imitation Policies
-Both DP3 ze2024dp3 and EquiBot yang2024equibot (refer Fig.
-12) ingest segmented 3D object point clouds but differ in network architectures.
+Figure 17: Qualitative real robot rollout samples for two tasks (plugpen and reorient) on the auxiliary dual-arm platform.
+IX-B Implementation of Imitation Policies
+Both DP3 [98] and EquiBot [94] (refer Fig.
+13) ingest segmented 3D object point clouds but differ in network architectures.
 Specifically, DP3 employs specially designed multiple MLPs to encode point clouds into latent features, then directly regressing dual-arm keyposes.
 EquiBot enhances robustness via a SIM(3)-equivariant PointNet variant: point clouds undergo rotation/scale-invariant feature extraction before action prediction.
-This equips EquiBot with improved generalization to unseen object variations (refer Tab.
-1 in the main paper), achieving 5–12% higher success rates than DP3 under identical training data and settings.
+This equips EquiBot with improved generalization to unseen object variations, achieving 5–12% higher success rates than DP3 under identical training data and settings.
 Both policies operate in open-loop, aligning with synthesized demonstration formats.
@@
 Network Architecture.
-For DP3, it is a variant of Diffusion Policy chi2023diffusion with a simpler point cloud encoder.
+For DP3, it is a variant of Diffusion Policy [16] with a simpler point cloud encoder.
 It also designs a two-layer MLP to encode robot proprioceptive states before concatenating with the observation representation.
-For EquiBot, we use a SIM(3)-equivariant PointNet++ yang2024equivact with 4 layers and hidden dimensionality 128 as the feature encoder.
+For EquiBot, we use a SIM(3)-equivariant PointNet++ [95] with 4 layers and hidden dimensionality 128 as the feature encoder.
 For the noise prediction network, we inherits hyperparameters from the original Diffusion Policy.
-Specifically, to optimize for inference speed in all experiments, we use the DDIM scheduler song2021denoising with 8 denoising steps, instead of the DDPM scheduler ho2020denoising which performs up to 100 denoising steps.
+Specifically, to optimize for inference speed in all experiments, we use the DDIM scheduler [79] with 8 denoising steps, instead of the DDPM scheduler [35] which performs up to 100 denoising steps.
 Sampling of Point Cloud.
@@
 For every evaluation in the real world, we run the policy in a randomly initialized placement of objects, and record the average success rate achieved by the policy.
-Figure 13: Qualitative real rollout samples from the left RGB camera for all evaluation scenarios. From top to bottom, they are defined six bimanual tasks. Best to view after zooming in.
 In addition, we have adopted the object-centric point cloud input.
@@
 This core design relies on the still rapidly developing capabilities of vision foundation models (VFMs).
-Here we leverage SOTA open vocabulary detection method Florence-2 xiao2024florence and segmentation method SAM2 ravi2024sam to automatically extract object masks and then filter out corresponding point clouds.
+Here we leverage SOTA open vocabulary detection method Florence-2 [89] and segmentation method SAM2 [74] to automatically extract object masks and then filter out corresponding point clouds.
 Despite this, occasionally we may fail to segment desired objects accurately, and in these special cases we will manually correct the masks.
@@
 Because we believe that the next generation of VFMs can alleviate these problems, or we can directly address them through domain adaptation, test-time adaptation, or adjusting input prompts.
-Appendix D More Real Robot Results Visualization and Analysis
-D.1 More Examples from Real Rollouts
-As shown in Fig.
-13, it supplements more qualitative results with additional real-robot execution visualizations, highlighting nuanced state transitions and task-specific challenges.
+X More Results Visualization and Analysis
+X-A More Examples from Real Rollouts
+Fig.
+14 supplements more qualitative results with additional real-robot execution visualizations, highlighting nuanced state transitions and task-specific challenges.
 These visualizations underscore BiDemoSyn’s ability to handle real-world complexities (such as mechanical tolerances and imperfect perception), while also exposing limitations in dynamic force modulation (e.g., over-pressing bottles or lids).
 Please refer our Supplemental Videos which provide frame-by-frame analyses of these executions, further dissecting critical phase and offering insights into policy decision-making under uncertainty.
-D.2 Statistics and Analysis of Failed Cases
-While policies trained with BiDemoSyn achieve high success rates, failure analysis reveals systematic challenges.
-Note that although the policy we trained is executed end-to-end (it establishes an implicit mapping from observations to action outputs), we can still align the analysis from the stage where its failure cases are located to find the core cause of the error.
-Finally, using the experimental results of EquiBot under out-of-distribution evaluations, we categorize failures into five types according to the task execution logic (mainly from the design module of BiDemoSyn):
-Figure 14: Typical failure cases.
-Left four cases: These include two cases from the plugpen task that incorrectly detected the marker body, and two cases from the reorient+unscrew task that incorrectly estimated the lying bottle’s orientation.
-Middle six cases: These include three cases from the pouring task where water was spilled or the mug cup was misgrabbed, and three cases from the unscrew task where the gripper failed to grab the bottle’s body or cap accurately (these are extra footages from a third-person perspective).
-Right four cases: These include two cases in the inserting task where the robotic arm’s end effector occasionally obscures the marker pen, and two cases in the reorient task where human interference causes the hand to partially obscure the spoon/shovel.
-Here, visual perceptual defects do not necessarily lead to task failure, demonstrating the robustness of the proposed orientation estimation algorithm.
-  • Visual Perception Errors (15%): Inaccurate object segmentation (e.g., mistaking the tabletop as part of the marker pen) leads to misaligned grasps.
-  • Orientation Estimation Errors (32%): Incorrect object pose estimation (e.g., misjudging a spoon’s concave direction) causes malformed trajectories (e.g., flipping to the wrong side).
-  • Alignment Failures (18%): Vision-guided initial frame misalignment (e.g., bottle nozzle offset by >>1cm) propagates errors to subsequent steps (e.g., spilling during pouring).
-  • Initial Grasp Failures (28%): Collisions during pre-grasp motions (e.g., gripper nudging objects) or unstable grasps (e.g., slippage in reorient) prevent task initiation.
-  • Error Propagation (7%): Cumulative errors from earlier stages compound into irreversible failures (e.g., slightly misaligned pluging between pen body and pen cap in plugpen).
-As can be seen, the orientation estimation and initial grasp failures dominate, reflecting two core challenges: (1) current pose estimators struggle with symmetric or textureless objects (e.g., metal spoon or shovel), and (2) gripper-centric path planning lacks fine-grained contact modeling (e.g., avoiding pre-touch collisions for irregular shapes).
-Addressing these requires advances in category-agnostic pose estimation and short-horizon contact optimization, which are critical directions for our future work.
-Some visual failure examples are shown in Fig.
-14, which provides a foundation for a full understanding of the BiDemoSyn system and its robust adaptation to allow for dynamic pre-grasping.
-More dynamic rollouts can be found in our Supplementary Videos.
-D.3 Two More Complex, Long-Horizon Tasks
-To further evaluate the scalability of BiDemoSyn beyond the six primary bimanual tasks, we introduce two additional long-horizon tasks that require sequential manipulation steps and accumulate significantly more physical error throughout the execution:
+X-B Two More Complex Long-Horizon Tasks
+To further evaluate the scalability of BiDemoSyn beyond the introduced six bimanual tasks, we add two new long-horizon tasks that require sequential manipulation steps and accumulate more physical error throughout the execution:
   • reorient+unscrew: The robot must first upright a lying down bottle (reorient phase), then perform a dual-arm unscrewing motion to remove the cap (unscrew phase).
@@
 This composition integrates two high-precision subtasks and requires stable arm coordination across multiple transitions.
-Figure 15: Qualitative real rollout samples from the left RGB camera for two newly added complex long-horizon bimanual tasks. Best to view after zooming in.
-Data Synthesis Using BiDemoSyn.
-These two newly added tasks were synthesized following the same three-stage pipeline described in the main text.
+To utilize BiDemoSyn, these two newly added tasks were synthesized following the same three-stage pipeline described in the main text.
 We briefly summarize the configurations: for reorient+unscrew, we collected initial observations for 4 bottle instances, each placed at 7 orientations across 36 grid cells, yielding 1008 synthesized demonstrations; for unscrew+pouring, we collected initial observations for 2 bottle and 2 cup instances, each placed on 18 grid cells, yielding 1296 synthesized demonstrations.
 Importantly, both tasks introduce longer temporal horizons, sequential coordination, and increased compounding error, thus providing a stronger stress test than the original six tasks.
-Policy Training and Performance.
-Using the synthesized datasets, we trained visuomotor policies for each long-horizon task based on the adapted EquiBot architecture (the strongest performer in Tab.
-1 of the main text).
+Then, using the synthesized datasets, we trained visuomotor policies for each long-horizon task based on the adapted EquiBot architecture.
 Across both tasks, the trained policies exhibited high robustness to spatial variation (different placements and bottle orientations), reliable execution of multi-step dual-arm coordination, and consistent task success under mild external perturbations introduced during pre-grasp interpolation.
-Fig.
-15 presents some qualitative results with additional real-robot execution visualizations of these two new tasks.
-And Fig.
-16 provides illustrative examples of continuous dynamic interference scenarios among tasks pouring and reorient+unscrew.
+In Fig.
+15, we present some qualitative results with additional real-robot execution visualizations of these two new tasks.
+And in Fig.
+16, we provide illustrative examples of continuous dynamic interference scenarios among tasks pouring and reorient+unscrew.
 Full dynamic demonstrations and rollouts of all eight tasks (previous 6 tasks and the new 2 tasks) are included in the Supplemental Videos.
 These results further validate BiDemoSyn’s ability to rapidly scale to new, long-horizon tasks with minimal additional engineering, and reinforce its suitability as a practical framework for synthesizing large-scale, real-world training data.
-Figure 16: Qualitative examples of applying dynamic interferences during the pre-grasping execution.
-From top to bottom, they are segments of the dynamic closed-loop grasping phase of tasks pouring and reorient+unscrew, where each object is manually disturbed from one to three times.
+Figure 18: Qualitative examples of applying dynamic interferences in pre-grasping for the task reorient on the auxiliary dual-arm platform, where the object is manually disturbed up to five times.
 The red arrow indicates the direction of the manually moved object (interfering).
-The cyan arrow and yellow arrow indicate the movement direction of the left and right robotic arms (chasing) respectively.
-Best to view after zooming in.
-Table 2: Quantitative comparison results of success rates for the few-shot BiDemoSyn under in-distribution and out-of-distribution evaluations.
-The # means the BiDemoSyn training data size.
-Although the number of demonstrations used for training is always the same for each setting, the number of seed demonstrations used for trajectory synthesis is different.
-  in-distribution evaluations out-of-distribution evaluations
-Trained
-Policy
-Few-Shot
-BiDemoSyn
-plugpen
-inserting
-unscrew
-pouring
-pressing
-reorient
-Average
-Success
-Rate
-plugpen
-inserting
-unscrew
-pouring
-pressing
-reorient
-Average
-Success
-Rate
-#3888 #7776 #1152 #2592 #1296 #1008 #2916 #3888 #1008 #972 #324 #756
-  EquiBot 1-shot 28/30 28/30 26/30 25/30 24/30 25/30 86.7% 18/30 20/30 24/30 21/30 17/30 20/30 66.7%
-5-shot 28/30 28/30 26/30 26/30 25/30 25/30 87.8% 20/30 21/30 24/30 23/30 18/30 21/30 70.0%
-10-shot 28/30 28/30 26/30 27/30 25/30 26/30 88.9% 21/30 21/30 25/30 23/30 18/30 22/30 71.7%
-15-shot 28/30 28/30 27/30 27/30 26/30 26/30 90.0% 21/30 22/30 25/30 23/30 18/30 22/30 72.2%
-20-shot 28/30 28/30 27/30 27/30 26/30 26/30 90.0% 21/30 22/30 25/30 23/30 18/30 22/30 72.2%
-D.4 Few-Shot-Based Demonstration Synthesis
-While BiDemoSyn is designed around a one-shot paradigm, the framework naturally extends to a few-shot setting, where multiple real demonstrations may serve as complementary seeds.
-This subsection studies the effect of incorporating a small number of additional demonstrations (5–20 samples) obtained through standard teleoperation or kinesthetic teaching (used in this study).
-Few-Shot Integration into BiDemoSyn.
-Given multiple seed demonstrations, we incorporate them in two ways: (1) Enhanced Object Variation Coverage.
-Additional demonstrations expand the diversity of observed object geometries, grasp points, and intermediate contact poses.
-These are incorporated into the variant blocks during trajectory adaptation.
-(2) Improved Robustness of Reference Motion.
-Multiple demonstrations allow BiDemoSyn to aggregate or select more reliable canonical motion patterns for the invariant blocks, mitigating single-demonstration idiosyncrasies.
-Under these two settings, the followed overall synthesis pipeline remains unchanged.
-Experimental Findings.
-We have summarized results of this ablation study in Tab.
-2.
-These new experiments show: (1) In-distribution (ID) performance.
-Adding few-shot data provides little to no improvement.
-Policies trained with BiDemoSyn’s one-shot–generated datasets already exhibit near-saturation performance for seen objects and placements.
-(2) Out-of-distribution (OOD) performance.
-Few-shot data provides moderate gains, primarily by increasing the number of observed object instances and geometric variations, thereby improving generalization to unseen shapes and dimensions.
-These findings align with the intuition that BiDemoSyn already produces high-quality, physically faithful demonstrations from a single demonstration (refer Fig.
-5).
-Additional few-shot demonstrations mainly help by expanding the representational coverage rather than correcting failure cases caused by hardware limitations of the embodied robot entity itself.
-Additional Conclusion.
-The few-shot extension confirms that BiDemoSyn supports scalable integration of multiple demonstrations, maintains strong performance in the one-shot setting, benefits from few-shot seeds mostly in the OOD regime, and requires no modification to its core synthesis pipeline.
-This suggests a promising avenue for future research: exploring hybrid one-shot + few-shot paradigms that balance data-collection cost with maximal generalization capability.
-D.5 VFM-based Object Anchoring in Clutter
-To evaluate the applicability of BiDemoSyn under more realistic and visually complex environments, we additionally examine its object-anchoring capability when the target objects are placed in cluttered scenes with distractors, varied background patterns, and partial occlusions.
-Although the main experiments intentionally isolate environmental variables to study the core challenge of real-world demonstration synthesis, BiDemoSyn is fully compatible with more complex perception conditions due to its reliance on Vision Foundation Models (VFMs) for object segmentation.
-Clutter-Aware Object Localization.
-We employ Florence-2 xiao2024florence and SAM2 ravi2024sam to detect and segment the task-relevant items (e.g., bottle, mug) using either text queries or exemplar-based prompts.
-In cluttered tabletop scenes, the VFM stack consistently produces high-quality masks, even when distractors share similar colors or shapes.
-Qualitatively, the predicted masks remain stable across moderate occlusions, enabling the system to isolate the target object with minimal ambiguity.
+The yellow arrow indicates the movement direction of the right robotic arm (chasing).
+X-C Cross-Embodiment Transferability Testing
+Beyond training and evaluating policies on the primary dual-arm platform (Fig.
+8 top), we further investigate the cross-embodiment transferability of policies trained with BiDemoSyn.
+Specifically, we deploy the learned visuomotor policies on a distinct humanoid-style dual-arm robot with different kinematic structures, arm mounting configurations, and workspace geometry (Fig.
+8 bottom).
+Importantly, no additional demonstrations, fine-tuning, or embodiment-specific retraining are introduced.
+We evaluate two representative bimanual tasks—plugpen and reorient—which respectively emphasize precise dual-arm alignment and large-range end-effector rotation.
+As illustrated in Fig.
+17, policies trained on the original platform can be directly executed on the new robot after a simple coordinate transformation of the predicted 6-DoF end-effector poses, without requiring explicit calibration of dynamic or contact parameters.
+Across extensive trials, both tasks achieve stable and consistent execution performance, comparable to those reported on the source platform.
+To further assess robustness, we introduce external perturbations during the pre-grasp stage (similar to Fig.
+16), including manual displacement and rotation of target objects.
+As shown in Fig.
+18, the policy can reliably recover from such disturbances and successfully complete the task, demonstrating that the transferred policy preserves not only task semantics but also robustness to moderate environmental variations.
+These qualitative results highlight a key advantage of BiDemoSyn: by adopting object-centric observations and embodiment-agnostic action representations, the resulting policies exhibit strong zero-shot transferability across heterogeneous robot embodiments, suggesting promising scalability to broader robotic platforms.
+More rollouts on this humanoid platform are in our Supplemental Videos.
+Figure 19: (Top) Failure cases in the task plugpen that wrongly detected the marker body, and the task reorient+unscrew that incorrectly estimated the lying bottle’s orientation.
+(Bottom) Failure cases in the task inserting where the robotic arm’s end effector occasionally obscures the marker pen, and the task reorient where human interference causes the hand to partially obscure the manipulated spoon/shovel.
+Figure 20: Top: Example cluttered tabletop scenes used to evaluate the perception front-end of BiDemoSyn.
+Each scene includes the target objects for the pouring task—two bottles (text prompts are blue bottle and green bottle) and two mugs (text prompts blue cup and green cup)—surrounded by visually distracting household items.
+Two additional tool-like distractors (brush, dustpan) are intentionally introduced to support future extensions such as dual-arm sweeping tasks.
+Bottom: High-quality 2D object masks produced by Florence-2 + SAM2 for all queried categories.
+X-D Statistics and Analysis of Failed Cases
+While policies trained with BiDemoSyn achieve high success rates, failure analysis reveals systematic challenges.
+Note that although the policy we trained is executed end-to-end (it establishes an implicit mapping from observations to action outputs), we can still align the analysis from the stage where its failure cases are located to find the core cause of the error.
+Finally, using the experimental results of EquiBot under out-of-distribution evaluations, we categorize failures into five types according to the task execution logic (mainly from the design module of BiDemoSyn):
+(1) Visual Perception Errors (15%): Inaccurate object segmentation (e.g., mistaking the tabletop as part of the marker pen) leads to misaligned grasps.
+(2) Orientation Estimation Errors (32%): Incorrect object pose estimation (e.g., misjudging a spoon’s concave direction) causes malformed trajectories (e.g., flipping to the wrong side).
+(3) Alignment Failures (18%): Vision-guided initial frame misalignment (e.g., bottle nozzle offset by >>1cm) propagates errors to subsequent steps (e.g., spilling during pouring).
+(4) Initial Grasp Failures (28%): Collisions during pre-grasp motions (e.g., gripper nudging objects) or unstable grasps (e.g., slippage in reorient) prevent task initiation.
+(5) Error Propagation (7%): Cumulative errors from earlier stages compound into irreversible failures (e.g., slightly misaligned pluging between pen body and pen cap in plugpen).
+As can be seen, the orientation estimation and initial grasp failures dominate, reflecting two core challenges: (1) current pose estimators struggle with symmetric or textureless objects (e.g., metal spoon or shovel), and (2) gripper-centric path planning lacks fine-grained contact modeling (e.g., avoiding pre-touch collisions for irregular shapes).
+Addressing these requires advances in category-agnostic pose estimation and short-horizon contact optimization, which are critical directions for our future work.
+Some visual failure examples are shown in Fig.
+19, which provides a foundation for a full understanding of the BiDemoSyn system and its robust adaptation to allow for dynamic pre-grasping.
+More dynamic rollouts can be found in our Supplementary Videos.
+X-E VFM-based Object Anchoring in Clutter
+To evaluate the applicability of BiDemoSyn under more realistic and visually complex environments, we additionally examine its object-anchoring capability when target objects are placed in cluttered scenes with distractors and partial occlusions.
+Although the main experiments intentionally isolate environmental variables to study the core challenge of real-world demonstration synthesis, BiDemoSyn is fully compatible with more complex perception conditions due to its reliance on VFMs for object segmentation.
+We utilize Florence-2 [89] and SAM2 [74] to detect and segment task-relevant items using either text queries or exemplar-based prompts.
 Fig.
-17 illustrates representative examples where both bottle and mug are correctly segmented despite surrounding tools, utensils, packaging, and nonrigid paper balls.
-Implications for Demonstration Synthesis.
-Because BiDemoSyn’s variant-block adaptation depends only on (1) reliable 2D masks and (2) centroid- and principal-axis–based orientation inference, accurate object anchoring is the only requirement for extending the system to cluttered environments.
-Our experiments show that the VFM-based perception pipeline provides sufficiently robust segmentation to support (a) pose-aligned trajectory modulation and (b) collision-aware grasp adjustments, even when the background is visually nonuniform.
-Limitations and Future Directions.
-While VFMs significantly enhance robustness to visual complexity, extreme occlusions or heavy inter-object entanglement may still degrade performance.
+20 illustrates representative examples where both bottle and mug are correctly segmented out of surrounding tools, utensils, packaging, and nonrigid paper balls.
+In cluttered tabletop scenes, the VFM consistently produces high-quality masks, even when distractors share similar color similarity and semantic ambiguity.
+Qualitatively, the predicted masks remain stable across moderate occlusions, enabling the system to isolate the target object with minimal ambiguity and providing reliable anchors for subsequent pose inference and variant-block adaptation in BiDemoSyn.
+Because BiDemoSyn’s variant-block adaptation depends only on reliable 2D masks along with centroid- and principal-axis–based orientation inference, accurate object anchoring is the only requirement for extending the system to cluttered environments.
+Our experiments show that the VFM-based perception pipeline provides robust segmentation to support pose-aligned trajectory modulation and collision-aware grasp adjustments, even when the background is visually nonuniform.
+However, while VFMs significantly enhance robustness to visual complexity, extreme occlusions or heavy inter-object entanglement may still degrade performance.
 Extending BiDemoSyn to handle these cases will require multi-view sensing, active view planning, or integrating VLM-driven object-query disambiguation.
 Nevertheless, these initial results demonstrate that VFM-based anchoring provides a practical and reliable foundation for deploying BiDemoSyn beyond clean tabletop settings, thereby enabling future extensions to household, warehouse, and mobile manipulation scenarios.
-Figure 17: Top row: Example cluttered tabletop scenes used to evaluate the perception front-end of BiDemoSyn.
-Each scene includes the target objects for the pouring task—two bottles (text prompts are blue bottle and green bottle) and two mugs (text prompts blue cup and green cup)—surrounded by visually distracting household items.
-Two additional tool-like distractors (brush, dustpan) are intentionally introduced to support future extensions such as dual-arm sweeping tasks.
-Bottom row: High-quality 2D object masks produced by Florence-2 + SAM2 for all queried categories.
-Despite clutter, color similarity, and semantic ambiguity, the VFM stack consistently isolates each target object with precise boundaries, providing reliable anchors for subsequent pose inference and variant-block adaptation in BiDemoSyn.
-Appendix E Summary, Reflection and Outlook
-This work presents BiDemoSyn, a framework that bridges the reality gap in bimanual imitation learning by synthesizing diverse, physically feasible demonstrations from a single human example.
-Through systematic decomposition, vision-guided alignment, and hierarchical optimization, our method eliminates simulation dependencies while scaling data collection by orders of magnitude.
-Experiments across six contact-rich tasks validate its efficiency (5 seconds per demo), scalability, and generalization to unseen spatial and geometric variations.
-However, limitations persist: BiDemoSyn currently assumes quasi-static motions, struggles with highly dynamic tasks (e.g., catching), and faces challenges in extreme shape variations (e.g., deformable objects).
-Failures primarily stem from orientation estimation errors and grasp-time collisions, highlighting needs for category-agnostic pose estimators and contact-aware motion planners.
-Future work will extend BiDemoSyn to deformable object manipulation, integrate dynamic perception for moving targets, and explore hybrid learning paradigms (e.g., combining synthetic demos with reinforcement learning).
-By democratizing real-world demonstration synthesis, this work aims to accelerate progress toward generalizable robotic manipulation.
-Appendix F Use of Large Language Models
-In preparing this manuscript, we employed the GPT-4o language model strictly for text editing purposes, such as improving readability and polishing grammar.
-The model was not used to generate research ideas, design methodologies, conduct experiments, or analyze results.
-All substantive contributions remain entirely those of the authors.
+XI Summary, Reflection and Outlook
+This work introduces BiDemoSyn, a real-world demonstration synthesis framework that addresses a central bottleneck in bimanual imitation learning: how to scale high-quality, physically grounded data collection without reliance on simulation or extensive teleoperation.
+By decomposing a single kinesthetic demonstration into invariant coordination structures and object-dependent adaptations, and by coupling vision-guided alignment with lightweight trajectory optimization, BiDemoSyn enables the rapid generation of diverse, feasible bimanual demonstrations directly in the physical world.
+Extensive experiments across six contact-rich tasks demonstrate its efficiency, scalability, and strong generalization to novel object poses and geometries.
+Beyond the original one-shot setting, we further show that BiDemoSyn naturally extends to few-shot demonstration synthesis, where additional demonstrations primarily enhance out-of-distribution generalization without altering the core pipeline.
+Moreover, we demonstrate zero-shot cross-embodiment policy deployment, where policies trained on BiDemoSyn data transfer effectively to a distinct humanoid-style dual-arm robot, preserving high success rates across representative tasks.
+These results highlight the importance of object-centric observations and embodiment-agnostic action representations, and suggest that BiDemoSyn supports not only scalable data generation, but also robust policy reuse across hardware platforms.
+Despite these advances, several limitations remain.
+BiDemoSyn currently assumes quasi-static task execution and rigid or semi-rigid objects, making highly dynamic interactions and deformable object manipulation challenging.
+Failure cases are primarily attributed to perception noise—particularly orientation estimation for small or symmetric objects—and occasional contact inaccuracies during grasping and insertion.
+Addressing these issues will likely require tighter perception–control coupling, category-agnostic orientation reasoning, and more expressive contact-aware planners.
+Looking forward, we see multiple promising directions.
+Future work will explore extending BiDemoSyn to deformable and articulated objects, incorporating dynamic and multi-view perception, and integrating closed-loop feedback into the synthesis pipeline.
+More broadly, we envision BiDemoSyn as a foundation for hybrid data-centric learning paradigms, combining one-shot, few-shot, and synthesized demonstrations to balance data efficiency with maximal generalization.
+By lowering the barrier to acquiring large-scale, real-world bimanual data, this work aims to accelerate progress toward practical, general-purpose robotic manipulation.